### v1

# Human-Activity AGV Quality Assessment: A Benchmark Dataset and an Objective Evaluation Metric

Abstract
AI-driven video generation techniques have made significant progress in recent years. However, AI-generated videos (AGVs) involving human activities often exhibit substantial visual and semantic distortions, hindering the practical application of video generation technologies in real-world scenarios. To address this challenge, we conduct a pioneering study on human activity AGV quality assessment, focusing on visual quality evaluation and the identification of semantic distortions. First, we construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, consisting of $3,200$ AGVs derived from $8$ popular text-to-video (T2V) models using $400$ text prompts that describe diverse human activities. We conduct a subjective study to evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify semantic issues of human body parts. Based on Human-AGVQA, we benchmark the performance of T2V models and analyze their strengths and weaknesses in generating different categories of human activities. Second, we develop an objective evaluation metric, named AI-Generated Human activity Video Quality metric (GHVQ), to automatically analyze the quality of human activity AGVs. GHVQ systematically extracts human-focused quality features, AI-generated content-aware quality features, and temporal continuity features, making it a comprehensive and explainable quality metric for human activity AGVs. The extensive experimental results show that GHVQ outperforms existing quality metrics on the Human-AGVQA dataset by a large margin, demonstrating its efficacy in assessing the quality of human activity AGVs. The Human-AGVQA dataset and GHVQ metric will be released in public at GitHub.
1 Introduction
Figure 1: Overview of this work. This study aims to address the distortions in human activity AGVs, with some typical distortions shown in the first row. To achieve this goal, we introduce the Human-AGVQA dataset, consisting of $3,200$ human activity AGVs with the quality labels, and the GHVQ metric to objectively compute the quality of human activity AGVs.
AI-driven video generation, particularly in the form of text-to-video (T2V) generation [9, 1], can automatically create visual realism videos according to text descriptions. This convenient and cost-effective method of video creation has significant potential for applications in entertainment, art, advertising, education, and various other fields. However, many studies [33, 53] indicate that current T2V models still struggle to generate realistic human figures and actions. For instance, AI-generated videos (AGVs) may exhibit distorted, incomplete, or abnormal body parts and actions, as illustrated in Figure 1. These artifacts can significantly affect the user’s perceptual quality, as the human body is highly structured, and the human visual system is susceptible to irregular body form and movement irregularities.
Therefore, accurately assessing the quality of human-activity AGVs is critical to automatically monitor the visual quality of large-scale AGVs in video generation applications, measure the progress of T2V models, and serve as an optimization or reward function to enhance the capability of T2V models. Unfortunately, the quality assessment for AI-generated content (AIGC) [54] is still in its infancy. The general image/video quality assessment (I/VQA) metrics have been shown to perform poorly in evaluating AGVs [84, 39], while common used metrics in T2V studies, such as Inception Score (IS) [65], Fréchet Inception Distance (FID) [28], and Fréchet Video Distance (FVD) [79], evaluate the generation performance and diversity of T2V models by comparing a set of AGVs to real data distributions in feature Inception feature spaces, thus failing to reflect the quality at the individual video level.
To bridge this gap, we establish the first quality assessment benchmark for human activity AGVs, providing a comprehensive analysis of quality issues in human activity videos generated by mainstream T2V models and serving as a testbed for validating the effectiveness of quality assessment metrics in evaluating human activity AGVs. Specifically, we first construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, which contains $3,200$ AGVs derived by $8$ state-of-the-art T2V models using $400$ text prompts. We define two practical quality assessment problems for human activity AGVs. The first is quality scoring, which quantifies the specific quality score of AGVs, allowing us to compare the quality of one AGV or T2V model against others. The second one is semantic issue identification, which pinpoints problematic body parts that affect quality scores, providing insights for further optimization of T2V models. Therefore, for the quality scoring task, we invite $80$ subjects to rate the quality scores of AGVs based on three critical dimensions: human appearance quality, action continuity quality, and overall visual quality. For the semantic issue identification task, we invite $5$ experts to label whether the human bodies (face, arms, torso, legs, and feet) exhibit semantic issues. Based on these subjectively rated quality labels, we can benchmark the performance of T2V models and analyze their strengths and weaknesses in generating different categories of human activities.
To address the lack of objective metrics for assessing the quality of human activity AGVs, we propose the AI-Generated Human activity Video Quality metric (GHVQ) to automatically evaluate the quality scores of human activity AGVs and identifies their semantic issues. Specifically, GHVQ consists of a spatial quality analyzer, an action quality analyzer, a text feature extractor, and a quality regressor. The spatial quality analyzer extracts human-focused and holistic quality features at the frame level. For human-focused features, we employ body-part segment masks to explicitly extract human body part-aware features, followed by an inner-body distortion analysis module and a cross-body distortion analysis module to refine these features at the individual body part level and the body part interaction level, respectively. Holistic quality features are extracted using a pre-trained AIGC IQA method to better represent complex AIGC artifiacts. The action quality analyzer captures the temporal continuity of AGVs using a pre-trained action recognition model. To ensure semantic alignment, the text feature extractor is employed to capture text-based semantic features. Finally, the four types of features are concatenated to form quality-aware feature representations. Two multi-layer perceptrons (MLPs) are used to simultaneously predict quality scores and binary decisions for each assessment dimension. Experimental results demonstrate that GHVQ outperforms existing metrics across all evaluated quality dimensions on the Human-AGVQA dataset, highlighting its effectiveness as a comprehensive metric for assessing the quality of complex human activity AGVs. We summarize the contributions of this paper as:
  • We establish the Human-AGVQA dataset for assessing the quality of human activity AGVs. Human-AGVQA comprises $3,200$ videos generated by $8$ T2V models using $400$ text prompts, along with subjectively rated quality labels for $2$ types of assessment tasks. The diversity of text prompts and the richness of quality labels make our benchmark well-suited for investigating the quality assessment of human activity AGVs.
  • We propose the GHVQ metric to automatically evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify whether five human body parts exhibit semantic issues. Experimental results indicate that GHVQ significantly outperforms existing related quality metrics, demonstrating its effectiveness.
2 Related Works
Video Generation Techniques can be broadly categorized into three types: GAN/VAE-based methods [59, 62, 46, 14], autoregressive-based methods [86, 49, 31, 80, 3, 20], and diffusion-based methods [30, 25, 90, 104, 100, 8, 7, 67, 29, 35, 94]. Among these, diffusion-based methods have seen significant progress, with some commercial models [16, 1] already being developed or launched for real-world applications. However, several studies [70, 13, 53, 33] point out that current video generation models still struggle with generating realistic human characteristics and actions, which emphasizes the need for quality assessment study in human activity AGVs.
Video Quality Assessment studies can be divided into two categories: knowledge-driven methods and data-driven methods. Knowledge-driven BVQA methods leverage prior knowledge to extract quality features (e.g., natural scenes statistics features) for quality regression, such as TLVQM [38], VIDEVAL [77], RAPIQUE [78], etc. In contrast, data-driven BVQA methods automatically learn the quality-aware features by training a carefully designed deep neural network (DNN) in a learning-based manner. Popular data-driven methods generally 1) use a 2D network to extract key frame features and a 3D network to capture motion features from video chunks, as seen in models like BVQA [41], SimpleVQA [71], PatchVQA [96], MinimalisticVQA [73], etc., or 2) employ a 3D network directly to extract video-level features from video chunks, such as FastVQA [87]. The 2D and 3D backbones can be pre-trained on other computer vision tasks [95, 24] or fine-tuned on VQA datasets in an end-to-end manner. Since current VQA research primarily focuses on natural videos, these methods often show limited performance when assessing the quality of AGVs [52].
Figure 2: Illustration of the MOSs distribution. HA, AC, and OV quality represent human appearance, action continuity, and overall video, respectively. These abbreviations are used consistently throughout the paper.
AGV Quality Assessment has been explored in several studies [54, 33, 39]. For instance, FETV [53] benchmarks five T2V models based on static quality, temporal quality, overall alignment, and fine-grained alignment. EvalCrafter [54] evaluates 17 objective metrics, including visual quality, text-video alignment, motion quality, and temporal consistency, on a subjectively-rated AGV dataset. It also proposes a combined metric that leads to improved performance. VBench [33] introduces 16 additional dimensional metrics to evaluate the performance of T2V models. Moreover, [39] constructs the T2VQA-DB, a large-scale text-to-video quality assessment dataset containing 10,000 videos, and develops the T2VQA metric, which focuses on the overall quality of AGVs. Despite these advancements, no specific metric has been developed to evaluate the quality of human activity AGVs.
3 Human-AGVQA Dataset
3.1 Human activity AGVs Collection
Text Prompts Selection. Text prompts describe the video content generated by T2V models. To comprehensively analyze the quality assessment problem of human activity AGVs, the selected text prompts in our dataset should cover a wide range of real-world human activities. Therefore, we classify the properties of these words into ages, genders, ethnicity, occupations, scenes, emotion, appearance, and activities categories and further divide each category into $44$ subcategories that frequently occur in daily human life. We then randomly select one subcategory (e.g., sports) from the activity category, one subcategory (e.g., outdoor) from the scene category, and two or three subcategories from human appearance-based categories (e.g., athlete from the occupation category and male from the gender category). These keywords are combined into a text prompt using GPT-4 [61] with a prompt:
# Please construct a complete sentence based on the following four concepts: “sports”, “outdoor”, “athlete”, and “male”. Note that the specific words “sports”, “outdoor”, “athlete”, and “male” are not required to appear in the sentence, but the sentence should clearly convey the ideas related to these four categories. #.
We also use GPT-4 to evaluate the prompt’s validity, rejecting any that are deemed unreasonable. Using this method, we generate $400$ distinct text prompts describing various human activities. The specific subcategories along with their proportions are listed in the supplementary material.
Figure 3: The MOS comparison for $8$ T2V models.
T2V Models Selection. We select eight state-of-the-art (SOTA) T2V models, including Gen-2 [16], StableVideo [1], StreamingT2V [26], Video Crafter2 [9], MagicTime [98], Open Sora [40], Latte [55], and AnimateDiff [22], to generate the video for each prompt. We summary the detailed information about these T2V models in the supplementary material.
Figure 4: The MOS comparison for $8$ text categories and their $44$ subcategories.
In summary, a total of $3,200$ videos were generated by $8$ SOTA T2V models using $400$ diverse text prompts in the Human-AGVQA dataset.
3.2 Quality Assessment Dimensions
We evaluate the quality of human activity AGVs from two perspectives:
Visual Quality Scoring. We assess the quality of AGVs from three key dimensions: human appearance quality, action continuity quality, and overall video quality. Specifically, human appearance quality assesses whether the human bodies in the video are complete, semantically correct, and aligned with the prompt description. Action continuity quality focuses on the coherence of the person’s movements, the temporal consistency of limb actions, and whether the actions adhere to the laws of physical motion. Overall video quality evaluates the holistic video quality, including human appearance quality, action continuity quality, as well as background content quality. Participants rate each dimension on a scale from $1$ to $5$, where $1$ represents the lowest quality and $5$ represents the highest.
Semantic Distortion Identification. Most quality metrics only provide a specific numerical value to quantify video quality, and cannot indicate which parts of the video are distorted or need further enhancement or optimization. This limitation restricts their applicability in quality artifact analysis and fine-grained video quality optimization. To address this, we label the semantic distorted human body parts in AGVs, providing a detailed explanation of quality scores. Specifically, we first annotate whether five body parts (face, arms  [^1] Here, the arms include the hands., torso, legs, feet) are present in AGVs. Then, for each detected body part, we label whether it contains semantic artifacts. Finally, we provide $10$ binary labels per AGV for semantic distortion identification.
3.3 Subjective Quality Experiment
A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years. The group included $46$ males and $34$ females. Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification. $3,200$ videos were divided into eight groups, each containing $400$ videos that covered all 400 prompts, and each video was rated by $10$ subjects, and labeled by $5$ experts. In total, there are $96,000$ opinion scores and $160,000$ binary labels in the Human-AGVQA dataset. Additional details on the subjective experimental settings can be found in the supplemental material.
3.4 Data Processing and Analysis
For the quality scoring task, we follow the recommended method in  [2] to process the raw subjective ratings into the mean opinion scores (MOSs). For the distortion identification task, we use a voting method to determine body presence and body distortion. The label with the most votes is selected as the final result. The details about data processing and inter-subject consistency are shown in supplementary material.
3.4.1 MOS Distribution Analysis
Figure 2 shows the MOS distributions for the three quality dimensions, which follow a Gaussian distribution. This indicates that medium-quality AGVs outnumber both high- and low-quality AGVs. This suggests that the quality of human activity AGVs still needs improvement to meet visual quality standards. The average MOS for overall video quality is slightly higher than that of action continuity quality, and both are significantly higher than that of human appearance quality. This highlights that generating high-quality realistic humans remains a challenge for current T2V models, underscoring the importance of our study.
Figure 5: The percentage of AGVs that contain human body parts (a) and the percentage of these body parts exhibit semantic distortions (b).
Figure 6: The GHVQ framework comprises four essential modules: a spatial quality analyzer for extracting human-focused and holistic quality features, an action quality analyzer for capturing temporal continuity of the body movement, a text feature extractor for identifying video content that needs to be generated by text prompts, and a quality regressor for mapping these quality-aware features to multi-dimensional quality scores and semantic distortion labels.
3.4.2 MOS Analysis for T2V Models
We calculate the average MOSs of eight T2V models across three quality dimensions to assess their strengths and weaknesses in generating human activity videos. The results are shown in Figure 3. Gen-2 and StableVideo demonstrate superior performance compared to the other T2V models, especially in human appearance quality and overall video quality. Notably, both Gen-2 and StableVideo are commercial T2V tools, indicating that the generation quality of commercial algorithms continues to surpass that of academic open-source models. AnimateDiff and VideoCrafter2 show similar performance to Gen-2 and StableVideo in terms of action continuity quality, highlighting their temporal consistency modeling capabilities.
3.4.3 MOS Analysis for Text Prompts
Figure 4 shows the average MOS for $8$ text prompt categories and their $44$ subcategories. Firstly, for the categories related to human attributes such as age, gender, and race, there is no significant difference in the quality scores among their subcategories, indicating that the evaluated T2V models do not show bias toward human-related attributes. Secondly, video quality is generally lower in scenarios involving significant or frequent human movement, such as outdoor in the scene category, athletes and cleaners in the occupation category, and sports in the activity category. This may be because frequent body movements cause the degradation of human appearance. Thirdly, for the emotion category, simpler emotions such as sadness and happiness receive relatively high scores, whereas more complex emotions like anger, disgust, and anxiety are not as well-represented, suggesting that current T2V models continue to face challenges in accurately understanding and representing complex emotions. Regarding the appearance category, handsome and beautiful yield satisfactory results. That may be that the prompts like handsome and beautiful will make the generation model generate high-quality humans.
3.4.4 Semantic Distortion Identification
We counted all body parts appearing in the $3,200$ AGVs of the Human-AGVQA dataset and the distortions observed for each body part, as shown in Figure 5.
In Figure 5(a), we observe that the face, arms, and torso account for the largest proportions, while the legs and feet are significantly less than three parts. Notably, a small portion of the videos lack any visible body parts, indicating that some T2V models struggle to comprehend the text prompts and generate the corresponding human figures. In Figure 5(b), it is evident that the torso exhibits the fewest distortions among the five body parts, whereas the face, arms, and legs show more pronounced distortions. This suggests that current T2V models face significant challenges in generating complete and realistic body parts. The performance analysis of T2V models for distortion identification of human bodies can be found in the supplemental material.
4 Proposed Model
As illustrated in Figure 6, the proposed AI-generated human activity video quality (GHVQ) methods consists of four modules: a spatial quality analyzer, an action quality analyzer, a text feature extractor, and a quality regressor, which are detailed as follows:
4.1 Spatial Quality Analyzer
Assume a video $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{N-1}$ generated by a text prompt $p$, where each frame $\bm{x}_{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the $i$-th frame. Here, $H$ and $W$ are the height and width of each frame, and $N$ is the total number of frames. Given the substantial spatial redundancy between video frames, we first temporally downsample the video, $\bm{x}$, into a lower frame rate sequence, $\bm{y}=\{\bm{y}_{i}\}_{i=0}^{N_{s}-1}$, where $\bm{y}_{i}=\bm{x}_{\lfloor N/N_{s}\times i\rfloor}$, and $N_{s}$ denotes the total number of frames used for extracting spatial features.
We extract spatial quality features from two perspectives: human-focused and holistic, to more accurately evaluate the human appearance quality and overall video quality.
4.1.1 Human-focused Quality Feature Extraction
For frame ${\bm{y}}_{i}$, we first apply a human body-part segmentation model [36], to explicitly detect the body-part masks $M_{i}$, where $M_{i}\in\mathbb{R}^{C_{m}\times H\times W}$ represents five segmentation masks corresponding to the face, arms, torso, legs, feet. Then, we design a shallow hourglass network [60] to extract the high-resolution feature maps $F_{h,i}$ that preserve the resolution of human appearance while capturing rich low-level features. Here, $F_{h,i}\in\mathbb{R}^{C_{h}\times H\times W}$, with $C_{h}$ representing the number of channels of $F_{h,i}$.
To focus on features specific to human body regions, we multiply the feature maps $F_{h,i}$ with the body masks $M_{i}$ along the channel dimensions to derive the human body-aware feature maps $F_{{\rm body},i}$:
$F_{{\rm body},i}=F_{h,i}\cdot M_{i},$ (1)
where $F_{{\rm body},i}\in\mathbb{R}^{C_{m}\times C_{b}\times H\times W}$. Next, we develop two modules—the inner-body distortion analysis module (Inner-body DAM) and the cross-body distortion analysis module (Cross-body DAM)—to further extract human-focused quality features at the individual body part level and the body part interaction level, respectively.
Inner-body DAM. The inner-body DAM module is designed to capture the quality features related to each individual body part region. This allows the module to analyze each body part separately, helping to assess whether these body parts are presented in the video and which body parts exhibit quality issues. To achieve this, we aggregate the feature maps $F_{{\rm body},i}$ across the channel dimension $C_{h}$:
$F_{{\rm inner},i}=\sum_{j=0}^{C_{h}-1}F_{{\rm body},i}[,j:,:,:],$ (2)
where $F_{{\rm inner},i}\in\mathbb{R}^{C_{m}\times H\times W}$ contains $C_{m}$ feature maps, each focusing on quality analysis for one of the $C_{m}$ body parts.
Cross-body DAM. In contrast, the cross-body DAM module focuses on capturing interactive relationships between the body parts. Since a complete human action involves the coordinated movement of multiple body parts, jointly considering these interactions can better represent the quality of the human action. To achieve this, we aggregate the feature maps $F_{{\rm body},i}$ across the channel dimension $C_{m}$:
$F_{{\rm cross},i}=\sum_{j=0}^{C_{m}-1}F_{{\rm body},i}[j,:,:,:],$ (3)
where $F_{{\rm cross},i}\in\mathbb{R}^{C_{s}\times H\times W}$ contains $C_{s}$ feature maps focused on the quality analysis of all body parts.
Feature Refinement and Fusion. For $F_{{\rm inner},i}$ and $F_{{\rm cross},i}$, we apply a CNN followed by a Transformer encoder [23] to further refine the inner-body and cross-body features:
$F^{\prime}_{{\rm inner},i}={\rm Transformer}({\rm Conv}(F_{{\rm inner},i})),$ (4)
$F^{\prime}_{{\rm cross},i}={\rm Transformer}({\rm Conv}(F_{{\rm cross},i})),$
where $\rm Transformer$ denotes the Transformer encoder, and the $\rm Conv$ represents a two-layer CNN. Finally, we concatenate them and then apply a CNN layer followed by a self-attention layer [66] to derive the human-focused quality features:
$F_{{\rm bq},i}$ $={\rm Attention}({\rm Conv}([F^{\prime}_{{\rm inner},i},F^{\prime}_{{\rm cross},i}])),$ (5)
where $\rm Attention$ denotes the self-attention operator and $F_{{\rm bq},i}$ is the human-focused quality features.
4.1.2 Holistic Quality Feature Extraction
To extract quality-aware features of the entire frame, we pre-train a ViT [15] model on Pick-a-Pic [37], a large-scale text-to-image (T2I) quality assessment dataset, to learn the quality-aware feature representation for AIGC images. Then, for frame ${\bm{y}}_{i}$, we employ the pre-trained ViT [15] as the feature extractor to compute the holistic quality features:
$F_{{\rm hq},i}={\rm{ViT}}({\bm{y}}_{i}).$ (6)
Table 1: The performance comparison of existing IQA-based, AQA-based, and VQA-based metrics with our proposed GHVQ metrics on the Human-AGVQA dataset. NSC refers to natural scene content.
Quality Dimension Methods Pre-training / Initialization Model Type Zero-shot Fine-tuning
SRCC PLCC SRCC PLCC
Human Appearance NIQE (ISPL, 2012) [58] NA (handcraft) NSC 0.234 0.273 0.564 0.583
BRISQUE (TIP, 2012) [57] NA (handcraft) NSC 0.261 0.314 0.593 0.629
HyperIQA (CVPR, 2020) [68] TID2013 [63] NSC 0.314 0.339 0.684 0.705
UNIQUE (TIP, 2021) [101] KonIQ-10K [32] NSC 0.243 0.280 0.671 0.695
MUSIQ (ICCV, 2021) [34] KonIQ-10K [32] NSC 0.229 0.262 0.584 0.609
StairIQA (JSTSP, 2023) [72] KonIQ-10K [32] NSC 0.334 0.393 0.665 0.671
CLIP-IQA (AAAI, 2023) [81] KonIQ-10K [32] NSC 0.329 0.368 0.672 0.683
LIQE (CVPR, 2023) [102] KonIQ-10k [32] NSC 0.285 0.332 0.641 0.664
MA-AGIQA (ACMMM, 2024) [84] AGIQA-3k [43] AIGC 0.377 0.380 0.728 0.735
Q-Align (ICML, 2024) [89] fused  [32][17][96][21] NSC 0.364 0.419
GHVQ (proposed) AIGC —- 0.812 0.821
Action Continuity ACTION-NET (ACMMM, 2020) [99] Kinetics [6] NSC 0.197 0.221 0.541 0.551
USDL (CVPR, 2020) [74] Kinetics [6] NSC 0.207 0.260 0.583 0.582
CoRe (ICCV, 2021) [97] Kinetics [6] NSC 0.179 0.213 0.560 0.561
TSA (CVPR, 2022) [92] Kinetics [6] NSC 0.202 0.258 0.607 0.611
Motion Smoothness (CVPR, 2024) [33] AMT [48] AIGC 0.249 0.30
Temporal Flickering (CVPR, 2024) [33] RAFT [75] AIGC 0.109 0.238
Action-Score (CVPR, 2024) [54] VideoMAE V2 [82] AIGC 0.208 0.284
Flow-Score (CVPR, 2024) [54] RAFT [75] AIGC 0.153 0.255
GHVQ (proposed) AIGC 0.788 0.792
Overall Video TLVQM (TIP, 2019) [38] NA (handcraft) NSC 0.275 0.314 0.613 0.624
RAPIQUE (JSP, 2021) [78] NA (handcraft) NSC 0.313 0.351 0.632 0.648
VIDEAL (TIP, 2021) [77] NA (handcraft) NSC 0.338 0.357 0.638 0.651
PatchVQ (CVPR, 2021) [96] LSVQ [96] NSC 0.378 0.404 0.667 0.703
SimpleVQA (ACMMM, 2022) [71] LSVQ [96] NSC 0.368 0.380 0.695 0.709
FastVQA (ECCV, 2023) [87] LSVQ [96] NSC 0.385 0.408 0.708 0.733
DOVER (ICCV, 2023) [88] DIVIDE [88] NSC 0.315 0.337 0.713 0.736
T2VQA (ACMMM, 2024) [39] T2VQA [39] AIGC 0.357 0.364 0.751 0.754
UGVQ (Arxiv, 2024) [103] LGVQ [103] AIGC 0.349 0.358 0.748 0.757
EvalCrafter (CVPR, 2024) [54] DIVIDE [88] NSC 0.328 0.341
Q-Align (ICML, 2024) [89] fused  [32][17][96][21] AIGC 0.427 0.485
GHVQ (proposed) AIGC 0.780 0.788
Finally, we aggregate the human-focused features and the holistic quality features as the spatial quality features, then average them across all sampled frames:
$F_{{\rm sq},i}$ $=[F_{{\rm bq},i},F_{{\rm hq},i}],$ (7)
$F_{{\rm sq}}$ $=\sum^{Ns-1}_{i=0}F_{{\rm sq},i},$
where $F_{{\rm sq},i}$ and $F_{{\rm sq}}$ represent the spatial quality features of frame ${\bm{y}}_{i}$ and video sequence ${\bm{y}}$.
4.2 Action Quality Analyzer
The action quality analyzer is designed to quantify temporal continuity of videos, serving as a complement to the spatial quality analyzer in assessing comprehensively the quality of AGVs. To address this, we utilize an action recognition model, as the action quality analyzer to capture the temporal continuity of human body movement. Since action recognition models are trained on large-scale, authentic-captured human action datasets, such as the Kinetics series [6], they can effectively extract motion representations of human bodies. Moreover, previous VQA studies [71] have also demonstrated that these features are useful for video quality evaluation. Specifically, for video $\bm{x}$ and the action recognition network $\rm{SlowFast}$ [19], we calculate the action quality features $F_{\rm aq}$ using all frames in $\bm{x}$.
4.3 Text Feature Extractor
Since the purpose of T2V models is to generate video content that aligns with the text prompt, it is necessary to measure the alignment between the video content and the text prompt. Following most T2I and T2V alignment evaluation studies [33, 84], we utilize the text encoder of CLIP [37] to extract semantic features $F_{\rm tq}$ of the text prompt $p$.
4.4 Quality Regressor
The quality regressor is used to assess the three quality scores and determine whether human bodies are present and if there exist semantic distortions. So, we first integrate the spatial quality features, the action quality features, and the text features into the final quality-aware features $F_{q}$:
$F_{q}=[F_{\rm sq},F_{\rm aq},F_{\rm tq}].$ (8)
Then, $F_{q}$ are fed into two multi-layer perceptron (MLP) to produce the quality scores $\hat{q}$ and distortion labels $\hat{b}$.
The loss function consists of the mean absolute error (MAE) loss, rank loss [39], and binary cross-entropy loss. The MAE loss, and rank loss are used to optimize the quality scoring task, and the binary cross-entropy loss is used to optimize the distortion identification task.
5 Experiments
5.1 Experiment Settings
Compared Quality Metrics. Since no specific metrics have been designed to evaluate human appearance quality, we compared $10$ IQA models to assess the quality of video frames as a proxy for human appearance. For action continuity quality and overall video quality, we select $8$ action quality assessment (AQA) and $11$ VQA methods as comparison respectively. For body parts detection and their distortion identification, there are also no specific metrics, we benchmark $9$ general video large multi-modality (LMM) model for this task. A detailed introduction to these methods can be found in the supplemental material.
Dataset Splits. We split the Human-AGVQA dataset into $70$% for training, $10$% for validation, and $20$% for test. For video quality scoring, the compared metrics are evaluated on the test set using two approaches: zero-shot testing and fine-tuning testing. For body parts detection and distortion identification, we directly test zero-shot performance of compared LMMs.
Evaluation Criteria. We use PLCC and SRCC to evaluate the performance of the test metrics. Additionally, we use classification accuracy to assess the performance of body part detection and distortion identification.
Other experimental settings including the training details of GHVQ can be found in the supplemental material.
5.2 Performance Comparison
5.2.1 Quality Scoring Task
Human Appearance Quality. We observe that all compared IQA methods perform poorly in evaluating human appearance quality in the zero-shot setting, which highlights that distortions in human appearance in AGVs represent entirely new artifacts that existing metrics cannot handle. Fine-tuning on the Human-AGVQA dataset improves the performance of data-driven methods, but their results remain suboptimal. This may be because these methods focus on learning feature representations from global images, rather than human-focused regions. The proposed GHVQ achieves the best performance, outperforming the second-best method by $0.812$ and $0.821$ in terms of SRCC and PLCC, demonstrating the effectiveness of its design for extracting human appearance-focused features.
Table 2: The accuracy of video LMM models in identifying the occurrences and distortions of body parts in the Human-AGVQA dataset.
Models Occurrences / Distortions
Face Arms Torso Legs Feet Avg
LLaMAVID (7B) [47] 91.9 / 56.6 82.9 / 62.1 81.8 / 46.5 37.2 / 30.7 25.1 / 41.5 63.8 / 47.7
VideoChatGPT (7B) [56] 73.9 / 71.7 79.7 / 76.2 75.2 / 37.6 38.8 / 23.0 29.2 / 19.4 59.4 / 45.6
VideoLLaMA2 (7B) [11] 22.5 / 24.9 14.1 / 14.9 41.2 / 61.2 61.9 / 75.8 70.4 / 78.3 42.2 / 51.0
VILA1.5 (7B) [51] 22.4 / 36.6 20.1 / 22.5 16.5 / 60.7 60.8 / 74.7 71.3 / 76.8 38.2 / 54.3
NeXT-Video (7B) [44] 62.5 / 37.6 56.8 / 22.5 78.6 / 60.9 58.2 / 75.0 65.3 / 77.6 64.3 / 54.7
OneVison (7B) [42] 61.9 / 29.8 71.9 / 19.8 64.6 / 62.2 52.2 / 77.0 67.2 / 80.3 63.6 / 53.8
Qwen2-VL (7B) [83] 63.1 / 44.3 57.9 / 25.7 71.0 / 60.4 62.6 / 76.1 74.1 / 79.9 70.3 / 57.3
GPT-4o  [61] 92.4 / 29.7 89.7 / 19.3 85.1 / 61.6 62.1 / 77.4 76.1 / 80.1 80.0 / 53.6
GPT-4o-mini  [61] 90.6 / 28.4 86.5 / 20.1 82.7 / 60.4 62.6 / 75.3 74.0 / 77.3 78.6 / 52.3
GHVQ (proposed) 92.6 / 72.3 87.1 / 80.6 87.1 / 63.4 63.5 / 78.1 77.2 / 81.5 81.5 / 75.2
Action Continuity Quality. Similar to human appearance quality, the compared AVA methods and temporal metrics of AIGC VQA methods show a very low correlation in assessing action continuity quality. This is because AVA methods are typically designed to quantify the completeness of professional sports performances, while the temporal metrics in AIGC VQA methods rely on handcrafted motion descriptors that do not capture semantic content, such as those involving individuals. In the fine-tuning results, GHVQ outperforms the re-trained AVA methods, surpassing the second-best method by $0.788$ and $0.792$ in terms of SRCC and PLCC, demonstrating that the proposed model architecture is better suited to measure action continuity quality.
Overall Video Quality. The performance of VQA metrics in assessing overall video quality is slightly superior to that of other metrics for human appearance and action continuity quality. Generally, we find that AIGC-based VQA methods outperform general data-driven VQA methods, and both of these are more effective than knowledge-driven VQA methods. Similarly, re-training these methods leads to significant improvements. Since GHVQ comprehensively extracts human-focused, action continuity, and holistic frame quality features, the proposed GHVQ outperforms existing VQA methods, improving performance by $0.780$ and $0.788$ in terms of SRCC and PLCC.
Table 3: Ablation study results for different modules across quality metrics. SQA, AQA, and TFE refer to the spatial quality analyzer, action quality analyzer, and text feature extractor, respectively. The results for the three quality scores are reported using SRCC/PLCC, while distortion identification is presented in terms of occurrence and distortion accuracy.
SQA AQA TFE HA Quality AC Quality OV Quality Distortion Identification
0.633 / 0.638 0.617 / 0.631 0.724 / 0.733 0.575 / 0.512
0.767 / 0.770 0.747 / 0.748 0.728 / 0.740 0.753 / 0.704
0.781 / 0.789 0.754 / 0.765 0.773 / 0.776 0.802 / 0.736
0.812 / 0.821 0.788 / 0.792 0.780 / 0.788 0.815 / 0.752
5.2.2 Body Presence and Distortion Identifying Task
Identifying the presence and distortions of human body parts is a novel task, and we benchmark the performance of video LMM models for this task in Table 2. As for LMMS, GPT-4 series perform best in terms of detecting the presence of body parts, while Qwen2-VL achieve the highest accuracy in identifying whether body parts exhibit semantic distortions. We observe that that identifying distortions in arms is the most challenging, followed by facial distortions while detecting distortions in legs and feet is relatively easier. This aligns with the inherent complexity of structures like faces and arms (particularly hands). Overall, our GHVQ model outperforms in both presence detection and distortion identification tasks, improving accuracy by $1.8\%$ and $28.3\%$, respectively.
5.3 Ablation Study
We conducted an ablation study to validate the effectiveness of the spatial quality analyzer, action quality analyzer, and text feature extractor. The experimental results are shown in Table 3. We observe that the performance across all three quality dimensions and distortion identification degrades the most when the spatial quality analyzer is removed, highlighting the importance of extracting human-focused and AIGC-aware quality features. The action quality analyzer is less crucial than the spatial quality analyzer, possibly because the pre-trained action recognition model is less sensitive to motion in AGVs. The text feature extractor has the least impact compared to the spatial and action quality analyzers, as it only captures video content based on text prompts and does not directly engage with AGVs. In summary, the combination of all three modules achieves the highest performance across all quality dimensions.
6 Conclusion
In this paper, we present an in-depth quality assessment study on human activity AGVs, which includes the construction of the large-scale human activity AGV dataset Human-AGVQA and the development of the objective quality metric GHVQ for human activity AGVs. The diversity of human activities and the richness of quality labels in Human-AGVQA make it well-suited for developing and validating quality metrics for human activity AGVs. The GHVQ metric has demonstrated strong performance in assessing these videos, making it a valuable tool for T2V studies to measure progress in generating human activity content. We hope these contributions will promote the development and application of T2V models.
Supplementary Material
In this supplementary file, we provide more details of Human-AGVQA dataset in Section 7, including text prompts selection, T2V model details, and subjective assessment experiments. We then provide data processing details and further analysis in Section 8, including MOS analysis for T2V models, distortion identification analysis, and the relationship between distortion identification and visual quality. The detailed experiment results are discussed in Section 9, including evaluation criteria in our experiments, compared quality metrics in the proposed benchmark, training details of proposed GHVQ metric, and ablation study of different backbones.
7 Human-AGVQA Dataset
7.1 Text Prompts Selection
Figure 7: The distribution of 44 subcategories in 8 categories of the 400 prompts.
To ensure that the text prompts in our dataset encompass a wide range of real-world human activities, we perform a comprehensive and systematic classification of the words used to create human activity-oriented text prompts. Specifically, we first divide the words into $8$ categories according to their properties: ages, genders, races, occupations, scenes, emotion, appearance, and activities. Each of these categories is then further subdivided into more specific subcategories that are commonly observed in daily human life, as illustrated in Table 4.
Some example text prompts and their corresponding generated videos are illustrated in Figure 9. The proportions of each subcategory within their respective categories are illustrated in Figure 7. It can be observed that the proportions of subcategories in the age, occupation, gender, and ethnicity categories are identical. While for other categories, the subcategories venue and outdoor in scene, sadness in emotion, robust in appearance, and work and socializing in activity have larger proportions than their respective subcategories. In addition, we analyzed the length of text prompts in the Human-AGVQA dataset, and the distribution of word counts are presented in Figure 8. The average text prompt length is $13.85$, with a median of $14$ words.
Figure 8: Histogram and density plot of word count per prompt.
7.2 T2V models
Detailed information about eight T2V methods is presented in Table 5. All selected T2V models have been published within the past two years and represent the most advanced video generation capabilities of T2V models to date. Except for VideoCrafter2, the resolution of videos generated by these models is at least $512$, indicating high clarity. However, the frame rates of StreamingT2V, Latte, and AnimateDiff are below $10$ fps, which may result in low motion consistency in actions.
7.3 Subjective Quality Assessment Experiments
7.3.1 Visual Quality Scoring Criteria
Table 4: Subcategories of ages, genders, races, occupations, scenes, emotion, appearance, and activities.
Category Subcategories
age child, youth, middle-aged, elderly
gender male, female
scene outdoor, residence, venue, wild, workplace
appearance tall, slim, short, handsome, robust, beautiful
occupation teacher, doctor, engineer, journalist, chef, painter, farmer, athlete, cleaner, student
emotion happiness, sadness, surprise, anger, fear, anxiety, excitement, disgust
ethnicity white race, yellow race, black race
activity study, work, socializing, entertainment, wilderness, sports
The quality scores of AGVs are rated from three dimensions: human appearance quality, action continuity quality, and overall video quality. Participants rate each dimension on a scale from $1$ to $5$, where $1$ represents the lowest quality and $5$ represents the highest. For each dimension, the detailed rating criteria are listed as follows:
– Human Appearance Quality
  • 5 (Excellent): The human appearance in the video is highly realistic, detailed, and perfectly matches the prompt description. Facial features, body proportions, and skin textures are sharp, clear, and lifelike. Clothing details are well-defined, with no visible distortions, blurring, or artifacts. The overall appearance is completely consistent with the prompt, including attributes like age, gender, or specific physical traits described.
  • 4 (Good): The human appearance is generally realistic and matches the majority of the prompt description. Facial details and body proportions are clear, but some finer details may be missing. Slight blurring or mild artifacts may occasionally appear but do not significantly detract from the realism. Minor deviations from the prompt description (e.g., subtle inconsistencies in age or clothing details) may be present but remain acceptable.
  • 3 (Fair): The human appearance shows noticeable flaws but still conveys the general idea described in the prompt. Facial features and body details are somewhat blurred or distorted, and artifacts appear more frequently. The appearance partially matches the prompt, but certain elements (e.g., age, clothing, or specific features) may be misrepresented or missing.
  • 2 (Poor): The human appearance is highly unrealistic or inconsistent with the prompt description. Facial features and body proportions are poorly defined, with frequent blurring, distortion, or unnatural textures. Key aspects of the prompt (e.g., age, gender, or specific traits) are misrepresented or missing entirely, reducing the video’s relevance and realism.
  • 1 (Bad): The human appearance is unrecognizable, highly distorted, or completely missing from the video, which is inconsistent with the prompt. Facial and body details are absent or severely flawed. If the video does not contain any human body parts (despite the prompt requiring them), this score is set to 1.
Table 5: Video formats generated by the eight T2V models in the Human-AGVQA dataset.
Methods Open Source Year Duration (s) FPS Resolution
Latte [55] 2023 $2.0$ $8$ $512\times 512$
AnimateDiff [22] 2023 $2.0$ $8$ $512\times 512$
Gen-2 [16] 2023 $4.0$ $24$ $1408\times 768$
VideoCrafter2 [9] 2024 $1.6$ $10$ $512\times 320$
MagicTime [98] 2024 $2.1$ $23$ $512\times 512$
StreamingT2V [26] 2024 $5.0$ $8$ $1280\times 720$
Open-sora-plan [40] 2024 $2.7$ $24$ $512\times 512$
StableVideo [1] 2024 $4.0$ $24$ $1024\times 576$
Figure 9: Some example frames of $8$ text categories and $8$ T2V models in proposed Human-AGVQA dataset.
Figure 10: Annotation Interface for Subjective Quality Experiment.
– Action Continuity Quality
  • 5 (Excellent): Actions are smooth, natural, and perfectly aligned with the prompt description throughout the video. Movements are logical, continuous, and free from any interruptions or jerks. The interaction between humans and the environment or objects (if applicable) is realistic and seamlessly integrated. The video fully reflects the activities described in the prompt.
  • 4 (Good): Actions are generally smooth and align well with the prompt, though minor inconsistencies or occasional interruptions in movement may be present. Slight jerks or subtle deviations from the described actions occur but do not disrupt the overall continuity or viewing experience.
  • 3 (Fair): Actions show noticeable flaws in continuity and may partially align with the prompt description. Jerky movements, illogical transitions, or interruptions are more frequent, reducing the fluidity of the motion. Some actions described in the prompt may be misrepresented or incomplete.
  • 2 (Poor): Actions are highly inconsistent or poorly aligned with the prompt. Frequent interruptions, jerks, or illogical transitions significantly disrupt the flow of movement. Actions may be missing key elements or appear unnatural in the context of the prompt.
  • 1 (Bad): Actions are disjointed, illogical, or completely unrelated to the prompt description. Movements are erratic and fail to convey the intended activity. If the video does not contain any human body parts (despite the prompt requiring them), this score is set to 1.
– Overall Video Quality
  • 5 (Excellent): The video quality is outstanding, with high sharpness, vibrant colors, stable playback, and well-balanced lighting throughout. The overall presentation is highly realistic, consistent, and engaging. The content perfectly matches the prompt in terms of both context and visual quality. There are no distortions or artifacts, and all details align with the intended description.
  • 4 (Good): The video quality is generally high, with minor imperfections such as slight blurring, small artifacts, or occasional lighting inconsistencies. The content mostly matches the prompt description, with only minor deviations or omissions. These issues do not significantly detract from the overall viewing experience.
  • 3 (Fair): The video quality is moderate, with noticeable flaws such as blurring, frequent artifacts, or inconsistent lighting. The prompt and video content partially align, but some key details may be missing or misrepresented. These issues reduce the overall coherence and impact of the video.
  • 2 (Poor): The video quality has significant issues, such as severe blurring, persistent artifacts, or poorly balanced lighting. The prompt and video content show weak alignment, with key elements missing or misinterpreted. These flaws make the video difficult to watch and significantly reduce its realism and coherence.
  • 1 (Bad): The video quality is extremely poor, with pervasive distortions, blurring, or lighting problems. The video content is almost entirely inconsistent with the prompt, failing to deliver the intended description. The video is barely watchable and lacks any sense of coherence.
7.3.2 Semantic Distortion Identification Criteria
– Human Body Presence
  • 1 (Present) A body part is labeled as present if its outline or general shape can be observed in the video, even if fine details are unclear due to distortion (e.g., blurry facial features or stiff limb movements). For example, the overall outline of the face is distinguishable, the arms are reasonably connected to the torso, and the shapes of the legs and feet are visible. Even minor proportional inconsistencies or blurriness do not disqualify a body part from being labeled as present.
  • 0 (Not Present) A body part is labeled as absent if it is entirely missing from the video (e.g., obscured, generation failure, or no relevant features are displayed) or if severe quality issues make its outline or shape unrecognizable (e.g., when the background and the body part are indistinguishable). For instance, if the face cannot be separated from the background, the arms or legs are entirely absent, or the feet are fused with the ground to an indistinguishable degree, the body part is considered absent.
– Human Body Distortion
  • 1 (Distorted) A body part is labeled as distorted if it shows significant quality issues. Examples include blurry facial features, disproportionate or distorted shapes, compression artifacts, arms with unnatural movements or disconnection from the torso, a torso with abnormal proportions or missing details, legs that appear broken, stretched, or disproportionate, and feet with unnatural shapes, incorrect positions, or unnatural interactions with the ground (e.g., floating or penetrating the surface).
  • 0 (No Distortion) A body part is labeled as no distortion if it appears clear, natural, and consistent with the overall video. For example, facial features are proportionate, arm movements are smooth and natural, the torso connects properly with other body parts, and the shapes and motions of the legs and feet align with expected behavior.
Figure 11: Annotation Interface for Distortion Identification.
7.3.3 Subjective Quality Experiment Procedure
A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years. The group included $46$ males and $34$ females. Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification. Each video was rated by $10$ subjects, and labeled by $5$ experts. The annotation interface for subjective quality experiment and distortion identification are shown in Figure 10 and Figure 11.
All subjects had normal or corrected-to-normal vision. The experiments were conducted in a controlled environment according to the recommendations of ITU-R BT.500-13 [2] to minimize external variables that could influence the judgments of the subjects. The setup included 27-inch calibrated display monitors with 95% DCI-P3 color gamut and a resolution of 4K. The viewing distance was set at $70$ cm. The room lighting was maintained at a consistent level of $300$ lux to ensure uniformity across all viewing sessions.
Before the formal assessments, subjects underwent a training session where they reviewed sample AGVs that were not included in the formal experiment. This session aimed to familiarize them with the evaluation criteria and the rating interface. In the formal experiment, $3,200$ videos were divided into eight groups, each containing $400$ videos that covered all 400 prompts. To avoid visual fatigue, each session lasted no longer than $30$ minutes, ensuring participants could maintain a high level of attention and accuracy in their ratings. In total, there are $96,000$ opinion scores and $160,000$ binary labels in the Human-AGVQA dataset.
8 Data Processing and Analysis
We follow the recommended method in  [2] to process the subjective ratings collected during the experiment. Outlier ratings are detected and removed if they deviate by more than $2\sigma$ (for normal distributions) or $\sqrt{20}\sigma$ (for non-normal distributions) from the mean rating for that condition. Observers contributing more than 5% of outlier ratings are excluded from the analysis. For each test condition, the mean score ($\mu_{i}$) and standard deviation ($\sigma_{i}$) are calculated based on all valid ratings provided by observer $i$:
$\mu_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}s_{ij},\quad\sigma_{i}=\sqrt{\frac{1}{N_{i}-1}\sum_{j=1}^{N_{i}}(s_{ij}-\mu_{i})^{2}}$ (9)
where $s_{ij}$ represents the raw rating assigned by observer $i$ to condition $j$, and $N_{i}$ is the number of conditions rated by observer $i$. To mitigate individual bias, each raw score $s_{ij}$ is normalized to a Z-score:
$Z_{ij}=\frac{s_{ij}-\mu_{i}}{\sigma_{i}}$ (10)
Finally, the Mean Opinion Score (MOS) for each test condition $j$ is computed as the average of the normalized Z-scores across all observers ($M_{j}$):
$MOS_{j}=\frac{1}{M_{j}}\sum_{i=1}^{M_{j}}Z_{ij}$ (11)
For distortion identification, we use a voting method to determine body presence and body distortion. The label with the highest number of votes is selected as the final result.
8.1 Inter-subject Consistency
To evaluate the consistency and reliability of quality scoring, we compute the inter-annotator agreement metric, Krippendorff’s Alpha ($\alpha$), where Krippendorff’s Alpha ($\alpha$) for HA quality, AC quality, OV quality are $0.708$, $0.695$, and $0.628$, respectively, indicating appropriate variations among subjects.
For semantic distortion identification, we calculated the proportions of options receiving 3 votes, 4 votes, and 5 votes, which were $9.7\%$, $41.5\%$, and $46.4\%$, respectively. This demonstrates that the majority of the votes from subjects were consistent.
8.2 MOS Analysis for Each T2V Model
We analyzed the MOS of three quality dimensions across $44$ attributes (mentioned in Table 4) for eight T2V models (mentioned in Table 5), as shown in Figure 12.
Figure 12: The MOS comparison for $8$ text categories of each T2V model.
StableVideo. Overall video (OV) quality emerges as the strongest aspect of StableVideo, consistently achieving high scores on most attributes, including complex scenarios such as workplace and emotion like anger. This highlights the model’s robust capability to generate visually coherent outputs. Human appearance (HA) quality also performs well, though it slightly lags behind overall video quality, indicating the model’s ability to generate clear and recognizable human features but with room for improvement compared to its overall visual performance. In contrast, action continuity (AC) quality represents the model’s primary weakness, consistently scoring the lowest among the three dimensions. This is particularly evident in subcategories requiring complex movements or high dynamism, such as surprise in emotion and sports in activity, revealing the model’s limitations in modeling motion continuity.
AnimateDiff. The AnimateDiff model excels in AC Quality, particularly in emotion and activity, where it consistently outperforms the other quality dimensions. This demonstrates the model’s strength in generating continuous and coherent motion. OV Quality shows stable and relatively high performance across most attributes, reflecting the model’s ability to maintain overall visual consistency. However, HA Quality falls behind, especially in categories such as occupation and activity, where the model struggles with rendering detailed and realistic human appearances. These findings suggest that while AnimateDiff is highly competent in motion continuity, it faces challenges in producing visually intricate details.
Gen-2. The Gen-2 model demonstrates exceptional performance in OV Quality, consistently achieving the highest scores across all subcategories. This underscores the model’s strength in generating visually cohesive outputs. HA Quality ranks second, showing a stable and strong ability to generate clear and recognizable human features, although it slightly trails OV Quality. AC Quality, however, is the weakest of the three dimensions, particularly in subcategories involving complex movements, such as angry in emotion and sports in activity, where it exhibits noticeable limitations.
Latte. Latte shows only average performance in terms of OV quality, but it is still slightly insufficient compared with other models. AC Quality exhibits greater variability, with strong performance in categories like sadness and surprise, but struggles in more complex occupational categories. HA Quality is the weakest dimension, particularly in occupation and activity categories, where the model demonstrates limitations in generating detailed and realistic human appearances. These observations suggest that the Latte model performs moderately well in maintaining overall video quality, its capacity to generate fine-grained human details requires further improvement.
MagicTime. MagicTime performs poorly in AC quality among the three quality scores, and there is a certain gap compared with other models. In addition, HA quality consistently underperforms, especially in categories involving complex emotions, movements, or appearances, highlighting the model’s limitations in rendering detailed human features. The model shows balanced performance in categories like scene and gender but exhibits significant variability in occupation and appearance, indicating its inconsistent generation quality depending on the attribute.
Open-Sora. The performance of OV Quality and AC Quality is not satisfactory, and they still fall short significantly compared to other models. In addition, HA quality consistently lags behind, particularly in the occupation, emotion, and activity categories, where the model struggles to render detailed and realistic human appearances. These results suggest that while Open-Sora excels in visual consistency and motion continuity, it faces challenges in human appearance generation.
StreamingT2V. StreamingT2V shows only average performance in terms of OV quality and HA quality, indicating its limited capability to generate sufficiently detailed human appearances. However, it falls short in AC quality, particularly in dynamic categories such as occupation (Athlete and Cleaner and emotion (surprise and anger), where the scores are significantly lower. These findings indicate that while StreamingT2V is effective in static scenarios, its handling of motion continuity in dynamic contexts requires further improvement.
VideoCrafter2. VideoCrafter2 shows balanced performance in OV quality and AC quality, with stable scores across most attributes. However, HA quality consistently underperforms, particularly in emotion categories such as anger and disgust, and activity categories such as sports and work, where it significantly trails OV and AC quality. These results highlight the model’s limitations in generating detailed and realistic human appearances. Nevertheless, the model demonstrates balanced performance in categories like age, scene, ethnicity, and gender, reflecting its comprehensive generation capability in these areas. However, its inconsistent performance in emotion, particularly in attributes like anger and disgust, underscores the challenges it faces in modeling complex emotional details.
8.3 Distortion Identification Analysis
Table 6: The total and proportion of occurrences and distortions for body parts in the $400$ videos generated by each model. Red represents the minimum value in a column.
Models Occurrences / Distortions
Face Arms Torso Legs Feet None / Total
Latte [55] 358 / 250 324 / 310 329 / 140 89 / 56 48 / 38 21 / 794
89.5 / 69.8 81.0 / 95.7 82.3 / 42.6 22.3 / 62.9 12.0 / 79.2 5.2 / 69.2
AnimateDiff [22] 373 / 245 352 / 316 357 / 88 183 / 67 108 / 63 22 / 779
93.2 / 65.7 88.0 / 89.8 89.2 / 24.6 45.8 / 36.6 27.0 / 58.3 5.5 / 56.7
Gen-2 [16] 383 / 250 357 / 329 361 / 134 138 / 86 93 / 72 10 / 871
95.8 / 65.3 89.2 / 92.2 90.2 / 37.1 34.5 / 62.3 23.3 / 77.4 2.5 / 65.4
StreamingT2V [26] 361 / 304 366 / 342 348 / 199 170 / 127 105 / 92 15 / 1064
90.2 / 84.2 91.5 / 93.4 87.0 / 57.2 42.5 / 74.7 26.3 / 87.6 3.7 / 78.8
VideoCrafter2 [9] 382 / 324 350 / 295 365 / 117 98 / 60 64 / 49 11 / 845
95.5 / 84.8 87.5 / 84.3 91.2 / 32.1 24.5 / 61.2 16.0 / 76.6 2.8 / 67.1
Open-sora-plan [40] 379 / 333 376 / 371 367 / 214 134 / 84 68 / 60 9 / 1062
94.8 / 87.9 94.0 / 98.7 91.7 / 58.3 33.5 / 62.7 17.0 / 88.2 2.2 / 80.2
MagicTime [98] 320 / 292 309 / 302 303 / 177 248 / 177 216 / 187 66 / 1135
80.0 / 91.2 77.2 / 97.7 75.7 / 58.4 62.0 / 71.4 54.0 / 86.6 16.5/ 81.3
StableVideo [1] 393 / 289 378 / 323 381 / 127 128 / 60 72 / 47 2 / 846
98.3 / 73.5 94.5 / 85.4 95.3 / 33.3 32.0 / 46.9 18.0 / 65.3 0.5 / 62.6
We calculate the proportion of occurrences and distortions for each body part of AGVs for different T2V models, as detailed in Table 6. For body occurrences, StableVideo performs the best, with the highest proportion of main body parts (face, arms, torso) present among these T2V models, and only $0.5\%$ of videos generated by StableVideo lack any human depiction. In contrast, MagicTime performs the worst, with $16.5\%$ of videos missing any human body parts. In terms of distortions, AnimateDiff and StableVideo exhibit better performance, with VideoCrafter2 and Gen-2 also showing comparable results. In contrast, Open-Sora and MagicTime show significant deficiencies, with distortions in certain body parts accounting for over $95$%. These findings closely align with the subjective quality scores of corresponding T2V models.
Moreover, an observation of significant concern is that $16.5\%$ of videos generated by MagicTime lack any depiction of human body parts, indicating a shortfall in the model’s ability to maintain consistency between textual descriptions and visual content. StableVideo demonstrates the most robust performance in this regard.
8.4 The Relationship between Semantic Distortion Identification and Visual Quality Scoring
Figure 13: The correlation between the semantic distortion identification and subjective quality.
We calculated the correlation between the three subjective quality scores and the distortion identification labels, as shown in Figure 13. Notably, when analyzing the relationship between body part distortions and quality scores, we included only videos in which the corresponding body part was present. For example, in assessing the correlation between facial distortion and quality scores, we considered only videos where a face was visible.
8.4.1 Body Parts Distortion
Overall, the quality of T2V models across the three evaluated dimensions demonstrates a certain degree of correlation with the distortions in the five body parts. Among these, facial distortion has the most significant impact on human appearance quality, followed by distortions in the torso and legs. This pattern aligns with the natural human perception process, which prioritizes facial recognition before focusing on the torso and other body parts.
Of the three quality dimensions, HA Quality exhibits the strongest correlation with body part distortions, followed by AC Quality, with OV Quality showing the weakest correlation. This suggests that body part distortions most directly influence perceptions of human appearance quality, with a lesser effect on the continuity of human actions, and the least impact on overall video quality.
Interestingly, for certain models, such as VideoCrafter2, quality scores show a negative correlation with distortions in the arms and feet. This implies that viewers’ attention is drawn more to prominent areas like the face and torso, making distortions in less visually critical body parts less influential.
8.4.2 Body Parts Presence
We also calculated the correlation between the presence of five body parts and the subjective quality scores. The results indicate that the presence of a face significantly impacts video quality, followed by the torso. This may be attributed to the design of the subjective experiment, where videos with no visible human parts were assigned the lowest HA and AC quality scores (1 point). Moreover, in most cases, if the face (head) is absent, other body parts are also missing.
For the MagicTime model, its quality scores show a strong correlation with the presence of a face, with the HA score reaching as high as $0.52$. This suggests that the presence of a face plays a crucial role in determining the quality scores for videos generated by MagicTime. Conversely, it also indicates that MagicTime frequently generates videos without human faces, consistent with our findings in Table 6.
Interestingly, the presence of legs and feet often correlates negatively with all three quality dimensions, suggesting that their inclusion tends to degrade overall video quality. This underscores the current limitations of T2V models in accurately generating human legs and feet, as distortions in these areas significantly impact viewers’ subjective perception of video quality across all dimensions.
9 Experiment Settings
9.1 Evaluation Metrics
Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC) are widely utilized metrics to quantify the relationship between predicted scores and ground truth scores. SRCC measures the monotonic relationship between two variables by comparing their rank orders. Unlike metrics that assume linearity, SRCC is robust to non-linear relationships, making it particularly well-suited for assessing relative rankings. It is mathematically defined as:
$SRCC=1-\frac{6\sum_{n=1}^{N}(v_{n}-p_{n})^{2}}{N(N^{2}-1)},$ (12)
where $N$ denotes the number of data points, $v_{n}$ represents the rank of the $n$-th ground truth score, and $p_{n}$ represents the rank of the corresponding predicted score. The value of SRCC ranges from 00 to $1$, with $1$ indicating a perfect monotonic increasing relationship, and 00 signifying the absence of any monotonic relationship. In the context of quality assessment, a high SRCC indicates that the predicted rankings are well-aligned with the rankings of ground truth scores, reflecting the model’s ability to capture the subjective perception of quality differences.
In contrast, the Pearson Linear Correlation Coefficient (PLCC) measures the strength of the linear correlation between predicted and ground truth scores, focusing on how closely the predicted values align with the true scores in absolute terms. It is expressed as:
$PLCC=\frac{\sum_{n=1}^{N}(y_{n}-\bar{y})(\hat{y}_{n}-\bar{\hat{y}})}{\sqrt{\sum_{n=1}^{N}(y_{n}-\bar{y})^{2}\sum_{n=1}^{N}(\hat{y}_{n}-\bar{\hat{y}})^{2}}},$ (13)
where $y_{n}$ and $\hat{y}_{n}$ denote the $n$-th ground truth and predicted scores, respectively, and $\bar{y}$ and $\bar{\hat{y}}$ are their corresponding means. Like SRCC, PLCC also ranges from 00 to $1$, where $1$ indicates a perfect positive linear correlation and 00 denotes no linear relationship. Unlike SRCC, which evaluates relative rankings, PLCC assesses the accuracy of predicted scores with respect to their numerical values, making it a critical metric for evaluating absolute predictive precision.
9.2 Compared Quality Metrics
9.2.1 Image Quality Assessment Methods
In all the IQA methods mentioned below, we uniformly extract 8 frames from each video in the Human-AGVQA dataset. These frames are then fed into the corresponding IQA methods to obtain 8 quality scores, and the average of these scores is taken as the method’s output.
NIQE [58] is a no-reference image quality assessment (IQA) model that evaluates the quality of images based on natural scene statistics (NSS) without relying on human-rated training data. For the Human-AGVQA dataset, we applied NIQE to assess image quality using selected video frames.
BRISQUE [57] leverages directional contrast features extracted from the Curvelet domain to assess the quality of contrast-distorted images. For the Human-AGVQA dataset, we utilized this approach to evaluate image quality based on its directional contrast features.
HyperIQA [68] introduces a self-adaptive hyper network for blind image quality assessment (BIQA) in the wild, addressing diverse content and distortion types in real-world images. For the Human-AGVQA dataset, we applied HyperIQA to assess image quality by adaptively learning perception rules and predicting quality scores.
UNIQUE [101] is a unified uncertainty-aware BIQA model designed to handle synthetic and realistic distortions. The model predicts both quality scores and associated uncertainties, enabling a probabilistic interpretation of image quality. For the Human-AGVQA dataset, we used UNIQUE to assess video frame quality based on its learning-to-rank framework.
MUSIQ [34] introduces a multi-scale image quality Transformer designed to handle full-resolution images with varying sizes and aspect ratios. The model uses hash-based 2D spatial embedding and scale embedding to capture positional and scale information effectively. For the Human-AGVQA dataset, we applied MUSIQ to evaluate video frame quality by leveraging its multi-scale feature extraction capability.
StairIQA [72] introduces a staircase network that integrates features from multiple convolutional layers of a CNN to capture better both low-level visual details and high-level semantic information for in-the-wild image quality assessment. For the Human-AGVQA dataset, we used the official implementation and adopted this approach to evaluate image quality without modifications.
CLIP-IQA [81] is a visual perception assessment model that utilizes the pre-trained CLIP framework to evaluate both the quality and abstract attributes of images. It introduces an antonym prompt pairing strategy (e.g., Good photo vs. Bad photo) to mitigate linguistic ambiguity and employs cosine similarity to predict image quality scores. For the Human-AGVQA dataset, we used the official implementation, applying CLIP-IQA to assess video frame quality.
LIQE [102] is a multitask learning framework for blind image quality assessment (BIQA) that leverages vision-language correspondence. LIQE uses CLIP for feature embeddings and optimizes tasks jointly by marginalizing probabilities over a textual template that describes scene, distortion, and quality. For the Human-AGVQA dataset, we used the official implementation, applying the quality assessment textual template to evaluate the video frame quality.
MA-AGIQA [84] introduces a framework for AI-generated image quality assessment by combining large multi-modality models (LMMs) with traditional deep neural networks (DNNs). For the Human-AGVQA dataset, we used the official implementation, applying MA-AGIQA to evaluate the video frame quality without modifications.
9.2.2 Video Quality Assessment Methods
TLVQM [38] is a two-level video quality model that first computes low-complexity features across the entire video and then selects representative frames based on these features to extract high-complexity features.
RAPIQUE [78] is a video quality assessment model that combines spatial-temporal scene statistics with high-level features extracted using deep CNNs. For the Human-AGVQA dataset, we used the official implementation, applying sparse frame sampling and integrating spatial and temporal features to evaluate video quality.
VIDEVAL [77] applies a feature selection strategy based on efficient blind VQA models. Using the official open-source code, we converted Human-AGVQA videos from RGB to YUV420 format for feature extraction.
Patch-VQ [96] is a video quality model that extracts spatial and temporal features from video frames and patches to analyze the relationship between local distortions and overall video quality. For the Human-AGVQA dataset, we used Patch-VQ’s official implementation, dividing each video into space-time patches for feature extraction and quality prediction.
SimpleVOA [71] uses an end-to-end network to extract spatial features directly from video frames and motion features to capture temporal distortions. A pre-trained SlowFast model processes 8 uniformly sampled frames rescaled to a height of 520. We used the official SimpleVOA model and finetuned it on Human-AGVQA.
FAST-VQA [87] introduces a grid mini-patch sampling (GMS) strategy to balance local quality assessment with global context while reducing computational costs for high-resolution videos. We used the official FAST-VQA-B model and finetuned it on Human-AGVQA.
DOVER [88] is a video quality evaluator that assesses quality from technical and aesthetic perspectives. We used the official code without modifications.
T2VQA [39] is a transformer-based model designed to evaluate text-to-video quality, focusing on text-video alignment and video fidelity. It uses BLIP and Swin-T for feature extraction, followed by feature fusion through a cross-attention module and quality regression using a large language model. On the Human-AGVQA dataset, we used the official T2VQA model and finetuned it on Human-AGVQA.
Table 7: Details of the model structure for the proposed model.
Training Hyper-Parameters Name/Value More Information
Action Quality Analyzer SlowFast-R50 parameter size: 34M
Holistic Quality Feature PickScore backbone: CLIP-ViT-H/14
Human body-part segmentation Spaiens parameter size: 2B
Human appearance feature shallow hourglass parameter size: 56M
Text Feature Extraction PickScore backbone: CLIP-ViT-H/14
UGVQ [103] is a unified framework for video quality assessment, focusing on spatial quality, temporal coherence, and text-to-video alignment. It utilizes features from CLIP for text and visual representations and SlowFast for motion representation. For the Human-AGVQA dataset, we employed UGVQ’s feature extraction and fusion modules to evaluate video quality comprehensively across these three dimensions.
EvalCrafter [54] is an evaluation framework for text-to-video (T2V) generative models, focusing on video quality, text-video alignment, motion quality, and temporal consistency. It uses a benchmark of 700 prompts and 17 evaluation metrics to analyze T2V model performance. For overall video quality, we directly used implementation code in EvalCrafter on the Human-AGVQA dataset without specific changes.
Q-Align [89] introduces a methodology for teaching large multi-modality models (LMMs) to assess video quality using discrete text-defined levels such as excellent, good, and poor. It converts mean opinion scores (MOS) into these rating levels for training and infers scores by weighting the predicted probabilities of the levels. For the Human-AGVQA dataset, we used the official code for inference without modifications.
9.2.3 Action Quality Assessment Methods
USDL [74] is an uncertainty-aware score distribution learning approach for action quality assessment (AQA), treating actions as instances associated with score distributions. For the Human-AGVQA dataset, since all videos contain more than 16 frames, each video is evenly divided into ten segments. Feature extraction is performed using an I3D backbone pre-trained on Kinetics. The final scores are normalized linearly, scaling the raw scores based on their minimum and maximum values to a range of [0, 100]. These normalized scores are then used to construct Gaussian distributions with the normalized value as the mean.
ACTION-NET [99] is a hybrid attention network tailored for action quality assessment (AQA) in long videos. It combines dynamic video information with static postures of action subjects in selected frames. For the dynamic stream, frames are sampled at a rate of 4 frames per second. For the static stream, the first, middle, and last frames are sampled, and the action subjects are cropped using the original detection algorithm.
Table 8: Details of the hyper-parameters for the model training.
Training Hyper-Parameters Name/Value
frame sampling for SQA $8$
frame resolution for SQA $512\times 512$
frame resolution for AQA $256\times 256$
batch size (videos) $16$
lr max $51e-$ 5
Ir schedule StepLR
decay ratio 0.9
decay interval 5
numerical precision float32
epoch $30$
optimizer Adam
CoRe [97] evaluates action quality by regressing relative scores against a reference video with similar attributes, such as action category. Due to differences in categorization strategies between the Human-AGVQA and AQA-7 datasets, we randomly selected an exemplar video generated by another T2V model within the same prompt category. Each video was divided into four equal segments, each containing four consecutive frames.
TSA [92] is a temporal segmentation attention module designed for procedure-aware cross-attention learning after spatial-temporal feature extraction. Similar to CoRe, each video is evenly divided into segments of four consecutive frames and processed using the I3D model.
For Motion Smoothness, Temporal Flickering, Action-Score, and Flow-Score metrics, we directly used their respective implementation code in VBench [33] and EvalCrafter [54] without specific changes.
9.3 Training details
The detailed structure of the proposed GHVQ metric is provided in Table 7. The corresponding hyperparameters for model training are listed in Table 8.
9.4 Ablation Study: The Effectiveness of Different Backbones
To further validate the effectiveness of the selected backbone networks, we conducted comparative experiments on four extraction modules: body-part segmentation network, holistic quality feature extraction network, text feature extraction network, and action quality analyzer.
9.4.1 Body-part Segmentation Network
For the body-part segmentation network, performance differences between methods are more pronounced, as shown in Table 9. Among them, Sapiens achieves the best results in HA and AC Quality. This superior performance can be attributed to Sapiens’ high accuracy in body part segmentation and pose estimation, allowing for precise identification of key body parts and thereby enhancing human appearance quality and action continuity. In contrast, earlier models such as JPPNet and Deeplabv3+ lack the segmentation precision and capability to handle complex poses, leading to weaker performance in overall quality metrics. Sapiens’ strong performance is also reflected in distortion identification.
Table 9: Ablation Study of Body Segmentation Network.
Methods HA Quality AC Quality OV Quality Distortion Identification
SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
None 0.659 0.671 0.692 0.704 0.745 0.752 0.581 0.524
OpenPose [4] 0.757 0.764 0.740 0.748 0.753 0.751 0.745 0.667
Deeplabv3+ [10] 0.712 0.725 0.724 0.736 0.758 0.767 0.649 0.546
JPPNet [50] 0.685 0.692 0.726 0.732 0.773 0.782 0.582 0.475
HRNet [69] 0.772 0.783 0.704 0.714 0.763 0.773 0.675 0.588
Sapiens [36] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
Table 10: Ablation Study of Text Feature Extraction Network.
Methods HA Quality AC Quality OV Quality Distortion Identification
SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
None 0.781 0.789 0.754 0.765 0.773 0.776 0.802 0.736
CLIPScore [27] 0.795 0.803 0.758 0.764 0.778 0.778 0.810 0.737
BLIP [45] 0.806 0.812 0.772 0.782 0.768 0.773 0.810 0.744
viCLIP [85] 0.803 0.807 0.769 0.773 0.767 0.774 0.813 0.739
ImageReward [93] 0.801 0.805 0.775 0.778 0.772 0.779 0.811 0.735
HPSv2 [91] 0.810 0.817 0.779 0.790 0.780 0.787 0.814 0.751
PickScore [37] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
9.4.2 Holistic Quality Feature Extraction Network
For holistic quality features, PickScore achieves the highest scores across all quality metrics, as shown in Table 11. PickScore effectively captures rich holistic quality features from the input video. Compared to other methods such as CLIPScore, ViCLIP, and CLIP (OpenAI), PickScore’s superior performance suggests that accurate holistic quality feature significantly benefits video quality evaluation. It is worth noting that the performance of HPS v2 and CLIP-H/14 (LAION) is very close to that of Pickscore.
Table 11: Ablation Study of Holistic Quality Feature.
Methods HA Quality AC Quality OV Quality Distortion Identification
SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
None 0.704 0.722 0.681 0.695 0.727 0.734 0.761 0.710
CLIPScore [27] 0.728 0.739 0.712 0.719 0.743 0.748 0.772 0.717
viCLIP [85] 0.723 0.731 0.723 0.733 0.732 0.744 0.789 0.731
CLIP-B/16 (OpenAI) [64] 0.726 0.735 0.717 0.728 0.738 0.749 0.792 0.735
CLIP-B/32 (OpenAI) [64] 0.739 0.745 0.731 0.742 0.746 0.752 0.787 0.729
CLIP-L/14 (OpenAI) [64] 0.752 0.757 0.743 0.749 0.751 0.766 0.798 0.740
CLIP-L/14 (LAION) [12] 0.776 0.785 0.768 0.777 0.760 0.768 0.802 0.743
CLIP-H/14 (LAION) [12] 0.799 0.812 0.785 0.791 0.772 0.780 0.811 0.745
CLIP-G/14 (LAION) [12] 0.795 0.805 0.782 0.786 0.768 0.778 0.805 0.737
HPSv2 [91] 0.809 0.814 0.788 0.790 0.779 0.784 0.814 0.750
PickScore [37] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
Table 12: Ablation Study of Action Quality Analyzer.
Methods HA Quality AC Quality OV Quality Distortion Identification
SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
None 0.767 0.770 0.747 0.748 0.728 0.740 0.753 0.704
I3D [5] 0.781 0.789 0.752 0.762 0.709 0.713 0.811 0.739
C3D [76] 0.797 0.798 0.760 0.773 0.709 0.713 0.811 0.739
X3D [18] 0.808 0.812 0.784 0.790 0.764 0.769 0.813 0.747
SlowFast [19] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
9.4.3 Text Feature Extraction Network
For text feature extraction, PickScore achieves the highest scores across all quality metrics, as shown in Table 10. PickScore effectively captures rich semantic information from the input text, providing a more comprehensive contextual understanding of the video content, which plays a critical role in enhancing overall quality assessment. Compared to other methods such as CLIPScore and ViCLIP, PickScore’s superior performance suggests that accurate semantic information significantly benefits video quality evaluation. It is worth noting that the performance of HPS v2 is very close to that of Pickscore. PickScore leverages the Pick-a-Pic dataset for training, which is specifically designed to optimize performance on preference comparison tasks. In contrast, HPS v2 is trained on the more extensive HPD v2 dataset, characterized by its larger scale and greater diversity in both image content and preference distributions.
9.4.4 Action Quality Analyzer
For the action quality analyzer, the SlowFast method demonstrates superior performance across all quality metrics, particularly in HA Quality and AC Quality, as shown in Table 12. In comparison, other methods such as I3D and C3D show slightly lower performance. The advantage of SlowFast lies in its ability to capture dynamic features across fast and slow motion, enabling comprehensive temporal information extraction. Additionally, SlowFast achieves the best results in distortion identification, indicating its strong capacity for detecting and recognizing distortions within video content.

### v2

@@
 Human-Activity AGV Quality Assessment: A Benchmark Dataset and an Objective Evaluation Metric
-Abstract
+Abstract.
 AI-driven video generation techniques have made significant progress in recent years.
@@
 To address this challenge, we conduct a pioneering study on human activity AGV quality assessment, focusing on visual quality evaluation and the identification of semantic distortions.
-First, we construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, consisting of $3,200$ AGVs derived from $8$ popular text-to-video (T2V) models using $400$ text prompts that describe diverse human activities.
+First, we construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, consisting of $6,000$ AGVs derived from $15$ popular text-to-video (T2V) models using $400$ text prompts that describe diverse human activities.
 We conduct a subjective study to evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify semantic issues of human body parts.
@@
 The extensive experimental results show that GHVQ outperforms existing quality metrics on the Human-AGVQA dataset by a large margin, demonstrating its efficacy in assessing the quality of human activity AGVs.
-The Human-AGVQA dataset and GHVQ metric will be released in public at GitHub.
-1 Introduction
-Figure 1: Overview of this work.
-This study aims to address the distortions in human activity AGVs, with some typical distortions shown in the first row.
-To achieve this goal, we introduce the Human-AGVQA dataset, consisting of $3,200$ human activity AGVs with the quality labels, and the GHVQ metric to objectively compute the quality of human activity AGVs.
-AI-driven video generation, particularly in the form of text-to-video (T2V) generation [9, 1], can automatically create visual realism videos according to text descriptions.
+The Human-AGVQA dataset and GHVQ metric will be released publicly.
+video quality assessment, AI-generated video, human activity video
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Computer vision problems
+Figure 1.
+Overview of this work.
+This study aims to address the distortions in human activity AGVs, with some typical distortions shown in the first column.
+To achieve this goal, we introduce the Human-AGVQA dataset, consisting of $6,000$ human activity AGVs with quality and distortion identification labels, and we benchmark over $40$ quality assessment and MLLM methods on it.
+Finally, the proposed GHVQ metric objectively computes the quality of human activity AGVs.
+1. Introduction
+AI-driven video generation, particularly in the form of text-to-video (T2V) generation (Chen et al., 2024a; AI, 2024a), can automatically create visual realism videos according to text descriptions.
 This convenient and cost-effective method of video creation has significant potential for applications in entertainment, art, advertising, education, and various other fields.
-However, many studies [33, 53] indicate that current T2V models still struggle to generate realistic human figures and actions.
+However, many studies (Huang et al., 2024; Liu et al., 2023) indicate that current T2V models still struggle to generate realistic human figures and actions.
 For instance, AI-generated videos (AGVs) may exhibit distorted, incomplete, or abnormal body parts and actions, as illustrated in Figure 1.
-These artifacts can significantly affect the user’s perceptual quality, as the human body is highly structured, and the human visual system is susceptible to irregular body form and movement irregularities.
+These artifacts can significantly affect the user’s perceptual quality, as the human body is highly structured, and the human visual system is susceptible to abnormal body parts and irregular movement.
 Therefore, accurately assessing the quality of human-activity AGVs is critical to automatically monitor the visual quality of large-scale AGVs in video generation applications, measure the progress of T2V models, and serve as an optimization or reward function to enhance the capability of T2V models.
-Unfortunately, the quality assessment for AI-generated content (AIGC) [54] is still in its infancy.
-The general image/video quality assessment (I/VQA) metrics have been shown to perform poorly in evaluating AGVs [84, 39], while common used metrics in T2V studies, such as Inception Score (IS) [65], Fréchet Inception Distance (FID) [28], and Fréchet Video Distance (FVD) [79], evaluate the generation performance and diversity of T2V models by comparing a set of AGVs to real data distributions in feature Inception feature spaces, thus failing to reflect the quality at the individual video level.
+Unfortunately, the quality assessment for AI-generated content (AIGC) (Liu et al., 2024a) is still in its infancy.
+The general image/video quality assessment (I/VQA) metrics have been shown to perform poorly in evaluating AGVs (Wang et al., 2024b; Kou et al., 2024), while common used metrics in T2V studies, such as Inception Score (IS) (Salimans et al., 2016), Fréchet Inception Distance (FID) (Heusel et al., 2017), and Fréchet Video Distance (FVD) (Unterthiner et al., 2018), evaluate the generation performance and diversity of T2V models by comparing a set of AGVs to real data distributions in feature Inception feature spaces, thus failing to reflect the quality at the individual video level.
 To bridge this gap, we establish the first quality assessment benchmark for human activity AGVs, providing a comprehensive analysis of quality issues in human activity videos generated by mainstream T2V models and serving as a testbed for validating the effectiveness of quality assessment metrics in evaluating human activity AGVs.
-Specifically, we first construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, which contains $3,200$ AGVs derived by $8$ state-of-the-art T2V models using $400$ text prompts.
+Specifically, we first construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, which contains $6,000$ AGVs derived by $15$ state-of-the-art T2V models using $400$ text prompts.
 We define two practical quality assessment problems for human activity AGVs.
@@
 Therefore, for the quality scoring task, we invite $80$ subjects to rate the quality scores of AGVs based on three critical dimensions: human appearance quality, action continuity quality, and overall visual quality.
-For the semantic issue identification task, we invite $5$ experts to label whether the human bodies (face, arms, torso, legs, and feet) exhibit semantic issues.
+For the semantic issue identification task, we invite $5$ experts to label whether the human bodies (face, hands, arms, torso, legs, and feet) exhibit semantic issues.
 Based on these subjectively rated quality labels, we can benchmark the performance of T2V models and analyze their strengths and weaknesses in generating different categories of human activities.
@@
   • We establish the Human-AGVQA dataset for assessing the quality of human activity AGVs.
-Human-AGVQA comprises $3,200$ videos generated by $8$ T2V models using $400$ text prompts, along with subjectively rated quality labels for $2$ types of assessment tasks.
+Human-AGVQA comprises $6,000$ videos generated by $15$ T2V models using $400$ text prompts, along with subjectively rated quality labels for $2$ types of assessment tasks.
 The diversity of text prompts and the richness of quality labels make our benchmark well-suited for investigating the quality assessment of human activity AGVs.
-  • We propose the GHVQ metric to automatically evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify whether five human body parts exhibit semantic issues.
+  • We propose the GHVQ metric to automatically evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify whether six human body parts exhibit semantic issues.
 Experimental results indicate that GHVQ significantly outperforms existing related quality metrics, demonstrating its effectiveness.
-2 Related Works
-Video Generation Techniques can be broadly categorized into three types: GAN/VAE-based methods [59, 62, 46, 14], autoregressive-based methods [86, 49, 31, 80, 3, 20], and diffusion-based methods [30, 25, 90, 104, 100, 8, 7, 67, 29, 35, 94].
-Among these, diffusion-based methods have seen significant progress, with some commercial models [16, 1] already being developed or launched for real-world applications.
-However, several studies [70, 13, 53, 33] point out that current video generation models still struggle with generating realistic human characteristics and actions, which emphasizes the need for quality assessment study in human activity AGVs.
-Video Quality Assessment studies can be divided into two categories: knowledge-driven methods and data-driven methods.
-Knowledge-driven BVQA methods leverage prior knowledge to extract quality features (e.g., natural scenes statistics features) for quality regression, such as TLVQM [38], VIDEVAL [77], RAPIQUE [78], etc.
-In contrast, data-driven BVQA methods automatically learn the quality-aware features by training a carefully designed deep neural network (DNN) in a learning-based manner.
-Popular data-driven methods generally 1) use a 2D network to extract key frame features and a 3D network to capture motion features from video chunks, as seen in models like BVQA [41], SimpleVQA [71], PatchVQA [96], MinimalisticVQA [73], etc., or 2) employ a 3D network directly to extract video-level features from video chunks, such as FastVQA [87].
-The 2D and 3D backbones can be pre-trained on other computer vision tasks [95, 24] or fine-tuned on VQA datasets in an end-to-end manner.
-Since current VQA research primarily focuses on natural videos, these methods often show limited performance when assessing the quality of AGVs [52].
-Figure 2: Illustration of the MOSs distribution.
+2. Related Works
+Video Generation Techniques can be broadly categorized into three types: GAN/VAE-based methods (Pan et al., 2017; Li et al., 2018), autoregressive-based methods (Liang et al., 2022; Hong et al., 2022; Villegas et al., 2022), and diffusion-based methods (Ho et al., 2022; Zeng et al., 2023; Chen et al., 2024b).
+Among these, diffusion-based methods have seen significant progress, with some commercial models (Esser et al., 2023; AI, 2024a) already being developed or launched for real-world applications.
+However, several studies (Sun et al., [n.
+d.]; Cho et al., 2024; Liu et al., 2023; Huang et al., 2024) point out that current video generation models still struggle with generating realistic human characteristics and actions, which emphasizes the need for quality assessment study in human activity AGVs.
+Figure 2.
+Illustration of the MOSs distribution.
 HA, AC, and OV quality represent human appearance, action continuity, and overall video, respectively.
 These abbreviations are used consistently throughout the paper.
-AGV Quality Assessment has been explored in several studies [54, 33, 39].
-For instance, FETV [53] benchmarks five T2V models based on static quality, temporal quality, overall alignment, and fine-grained alignment.
-EvalCrafter [54] evaluates 17 objective metrics, including visual quality, text-video alignment, motion quality, and temporal consistency, on a subjectively-rated AGV dataset.
+Video Quality Assessment studies can be divided into two categories: knowledge-driven methods and data-driven methods.
+Knowledge-driven VQA methods leverage prior knowledge to extract quality features (e.g., natural scenes statistics features) for quality regression, such as TLVQM (Korhonen, 2019), VIDEVAL (Tu et al., 2021a), RAPIQUE (Tu et al., 2021b), etc.
+In contrast, data-driven BVQA methods automatically learn the quality-aware features by training a carefully designed deep neural network (DNN).
+Popular data-driven VQA methods generally 1) use a 2D network to extract key frame features and a 3D network to capture motion features from video chunks, as seen in models like Li22 (Li et al., 2022b), SimpleVQA (Sun et al., 2022), PatchVQA (Ying et al., 2021), MinimalisticVQA (Sun et al., 2024), etc., or 2) employ a 3D network directly to extract video-level features from video chunks, such as FastVQA (Wu et al., 2022).
+The 2D and 3D backbones can be pre-trained on other computer vision tasks (Ying et al., 2020; He et al., 2016) or fine-tuned on VQA datasets in an end-to-end manner.
+Since current VQA research primarily focuses on natural videos, these methods often show limited performance when assessing the quality of AGVs (Liu et al., 2024b).
+AGV Quality Assessment has been explored in several studies (Liu et al., 2024a; Huang et al., 2024; Kou et al., 2024).
+For instance, FETV (Liu et al., 2023) benchmarks five T2V models based on static quality, temporal quality, overall alignment, and fine-grained alignment.
+EvalCrafter (Liu et al., 2024a) evaluates 17 objective metrics, including visual quality, text-video alignment, motion quality, and temporal consistency, on a subjectively-rated AGV dataset.
 It also proposes a combined metric that leads to improved performance.
-VBench [33] introduces 16 additional dimensional metrics to evaluate the performance of T2V models.
-Moreover, [39] constructs the T2VQA-DB, a large-scale text-to-video quality assessment dataset containing 10,000 videos, and develops the T2VQA metric, which focuses on the overall quality of AGVs.
+VBench (Huang et al., 2024) introduces 16 additional dimensional metrics to evaluate the performance of T2V models.
+Moreover, (Kou et al., 2024) constructs the T2VQA-DB, a large-scale text-to-video quality assessment dataset containing 10,000 videos, and develops the T2VQA metric, which focuses on the overall quality of AGVs.
 Despite these advancements, no specific metric has been developed to evaluate the quality of human activity AGVs.
-3 Human-AGVQA Dataset
-3.1 Human activity AGVs Collection
+Figure 3. The MOS comparison for $15$ T2V models.
+3. Human-AGVQA Dataset
+3.1. Human activity AGVs Collection
+Figure 4. The MOS comparison for $15$ text categories and their $44$ subcategories.
 Text Prompts Selection.
@@
 We then randomly select one subcategory (e.g., sports) from the activity category, one subcategory (e.g., outdoor) from the scene category, and two or three subcategories from human appearance-based categories (e.g., athlete from the occupation category and male from the gender category).
-These keywords are combined into a text prompt using GPT-4 [61] with a prompt:
-# Please construct a complete sentence based on the following four concepts: “sports”, “outdoor”, “athlete”, and “male”.
-Note that the specific words “sports”, “outdoor”, “athlete”, and “male” are not required to appear in the sentence, but the sentence should clearly convey the ideas related to these four categories.
+These keywords are combined into a text prompt using GPT-4 (OpenAI, 2024a) with a prompt:
+# Please construct a complete sentence based on the following four concepts: ”sports”, ”outdoor”, ”athlete”, and ”male”.
+Note that the specific words ”sports”, ”outdoor”, ”athlete”, and ”male” are not required to appear in the sentence, but the sentence should clearly convey the ideas related to these four categories.
 #.
@@
 The specific subcategories along with their proportions are listed in the supplementary material.
-Figure 3: The MOS comparison for $8$ T2V models.
 T2V Models Selection.
-We select eight state-of-the-art (SOTA) T2V models, including Gen-2 [16], StableVideo [1], StreamingT2V [26], Video Crafter2 [9], MagicTime [98], Open Sora [40], Latte [55], and AnimateDiff [22], to generate the video for each prompt.
+We select $15$ state-of-the-art (SOTA) T2V models, including Gen-2 (Esser et al., 2023), StableVideo (AI, 2024a), StreamingT2V (Henschel et al., 2024), Video Crafter2 (Chen et al., 2024a), MagicTime (Yuan et al., 2024), Open Sora (Lab and etc., 2024), Latte (Ma et al., 2024), and AnimateDiff (Guo et al., 2024), Sora(OpenAI, 2024b) , Dreamina(ByteDance, 2024), Kling(Technology, 2024), Ying(AI, 2024b), Wan(Cloud, 2025), LTX-Video(HaCohen et al., 2024), Show-1(Zhang et al., 2023a)to generate the video for each prompt.
 We summary the detailed information about these T2V models in the supplementary material.
-Figure 4: The MOS comparison for $8$ text categories and their $44$ subcategories.
-In summary, a total of $3,200$ videos were generated by $8$ SOTA T2V models using $400$ diverse text prompts in the Human-AGVQA dataset.
-3.2 Quality Assessment Dimensions
+In summary, a total of $6,000$ videos were generated by $15$ SOTA T2V models using $400$ diverse text prompts in the Human-AGVQA dataset.
+3.2. Quality Assessment Dimensions
 We evaluate the quality of human activity AGVs from two perspectives:
@@
 Participants rate each dimension on a scale from $1$ to $5$, where $1$ represents the lowest quality and $5$ represents the highest.
+Figure 5. The percentage of AGVs that contain human body parts (a) and the percentage of these body parts exhibit semantic distortions (b).
 Semantic Distortion Identification.
@@
 To address this, we label the semantic distorted human body parts in AGVs, providing a detailed explanation of quality scores.
-Specifically, we first annotate whether five body parts (face, arms  [^1] Here, the arms include the hands., torso, legs, feet) are present in AGVs.
+Specifically, we first annotate whether $6$ body parts (face, hands, arms, torso, legs, feet) are present in AGVs.
 Then, for each detected body part, we label whether it contains semantic artifacts.
-Finally, we provide $10$ binary labels per AGV for semantic distortion identification.
-3.3 Subjective Quality Experiment
+Finally, we provide $12$ binary labels per AGV for semantic distortion identification.
+3.3. Subjective Quality Experiment
 A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years.
@@
 Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification.
-$3,200$ videos were divided into eight groups, each containing $400$ videos that covered all 400 prompts, and each video was rated by $10$ subjects, and labeled by $5$ experts.
-In total, there are $96,000$ opinion scores and $160,000$ binary labels in the Human-AGVQA dataset.
+$6,000$ videos were divided into $15$ groups, each containing $400$ videos that covered all 400 prompts, and each video was rated by $15$ subjects, and labeled by $5$ experts.
+In total, there are $270,000$ opinion scores and $360,000$ binary labels in the Human-AGVQA dataset.
 Additional details on the subjective experimental settings can be found in the supplemental material.
-3.4 Data Processing and Analysis
-For the quality scoring task, we follow the recommended method in  [2] to process the raw subjective ratings into the mean opinion scores (MOSs).
+3.4. Data Processing and Analysis
+For the quality scoring task, we follow the recommended method in  (Met, 2002) to process the raw subjective ratings into the mean opinion scores (MOSs).
 For the distortion identification task, we use a voting method to determine body presence and body distortion.
@@
 The details about data processing and inter-subject consistency are shown in supplementary material.
-3.4.1 MOS Distribution Analysis
+3.4.1. MOS Distribution Analysis
 Figure 2 shows the MOS distributions for the three quality dimensions, which follow a Gaussian distribution.
@@
 This highlights that generating high-quality realistic humans remains a challenge for current T2V models, underscoring the importance of our study.
-Figure 5: The percentage of AGVs that contain human body parts (a) and the percentage of these body parts exhibit semantic distortions (b).
-Figure 6: The GHVQ framework comprises four essential modules: a spatial quality analyzer for extracting human-focused and holistic quality features, an action quality analyzer for capturing temporal continuity of the body movement, a text feature extractor for identifying video content that needs to be generated by text prompts, and a quality regressor for mapping these quality-aware features to multi-dimensional quality scores and semantic distortion labels.
-3.4.2 MOS Analysis for T2V Models
-We calculate the average MOSs of eight T2V models across three quality dimensions to assess their strengths and weaknesses in generating human activity videos.
+3.4.2. MOS Analysis for T2V Models
+We calculate the average MOSs of $15$ T2V models across three quality dimensions to assess their strengths and weaknesses in generating human activity videos.
 The results are shown in Figure 3.
-Gen-2 and StableVideo demonstrate superior performance compared to the other T2V models, especially in human appearance quality and overall video quality.
-Notably, both Gen-2 and StableVideo are commercial T2V tools, indicating that the generation quality of commercial algorithms continues to surpass that of academic open-source models.
-AnimateDiff and VideoCrafter2 show similar performance to Gen-2 and StableVideo in terms of action continuity quality, highlighting their temporal consistency modeling capabilities.
-3.4.3 MOS Analysis for Text Prompts
+Kling, Sora, and Wan are capable of producing highly realistic human appearances and action continuity with excellent overall quality, demonstrating strong generation quality.
+Dreamina, Ying, StableVideo, and Gen-2 also show strong performance.
+Notably, StableVideo and Gen2 exhibit significantly higher OV quality compared to the other two dimensions.
+These above models are well-known commercial T2V tools, reinforcing the trend that commercial algorithms often surpass open-source models in terms of output quality.
+LTX-Video, Show-1, AnimateDiff, and VideoCrafter2 show moderate performance across the dimensions, indicating that they have room for improvement.
+Latte, MagicTime, OpenSora, and StreamingT2V have lower scores across all three quality dimensions, which suggests that these models may face challenges in generating high-quality videos.
+3.4.3. MOS Analysis for Text Prompts
 Figure 4 shows the average MOS for $8$ text prompt categories and their $44$ subcategories.
@@
 That may be that the prompts like handsome and beautiful will make the generation model generate high-quality humans.
-3.4.4 Semantic Distortion Identification
-We counted all body parts appearing in the $3,200$ AGVs of the Human-AGVQA dataset and the distortions observed for each body part, as shown in Figure 5.
-In Figure 5(a), we observe that the face, arms, and torso account for the largest proportions, while the legs and feet are significantly less than three parts.
+Figure 6.
+The GHVQ framework comprises four essential modules: a spatial quality analyzer for extracting human-focused and holistic quality features, an action quality analyzer for capturing temporal continuity of the body movement, a text feature extractor for identifying video content that needs to be generated by text prompts, and a quality regressor for mapping these quality-aware features to multi-dimensional quality scores and semantic distortion labels.
+3.4.4. Semantic Distortion Identification
+We counted all body parts appearing in the $6,000$ AGVs of the Human-AGVQA dataset and the distortions observed for each body part, as shown in Figure 5.
+In Figure 5(a), we observe that the face, hands, arms, and torso account for the largest proportions, while the legs and feet are significantly less than four parts.
 Notably, a small portion of the videos lack any visible body parts, indicating that some T2V models struggle to comprehend the text prompts and generate the corresponding human figures.
-In Figure 5(b), it is evident that the torso exhibits the fewest distortions among the five body parts, whereas the face, arms, and legs show more pronounced distortions.
+In Figure 5(b), it is evident that the torso exhibits the fewest distortions among the six body parts, whereas the hands, face and feet show more pronounced distortions.
 This suggests that current T2V models face significant challenges in generating complete and realistic body parts.
 The performance analysis of T2V models for distortion identification of human bodies can be found in the supplemental material.
-4 Proposed Model
-As illustrated in Figure 6, the proposed AI-generated human activity video quality (GHVQ) methods consists of four modules: a spatial quality analyzer, an action quality analyzer, a text feature extractor, and a quality regressor, which are detailed as follows:
-4.1 Spatial Quality Analyzer
+4. Proposed Model
+As illustrated in Figure 6, the proposed AI-generated human activity video quality (GHVQ) method consists of four modules: a spatial quality analyzer, an action quality analyzer, a text feature extractor, and a quality regressor, which is detailed as follows:
+4.1. Spatial Quality Analyzer
 Assume a video $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{N-1}$ generated by a text prompt $p$, where each frame $\bm{x}_{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the $i$-th frame.
@@
 We extract spatial quality features from two perspectives: human-focused and holistic, to more accurately evaluate the human appearance quality and overall video quality.
-4.1.1 Human-focused Quality Feature Extraction
-For frame ${\bm{y}}_{i}$, we first apply a human body-part segmentation model [36], to explicitly detect the body-part masks $M_{i}$, where $M_{i}\in\mathbb{R}^{C_{m}\times H\times W}$ represents five segmentation masks corresponding to the face, arms, torso, legs, feet.
-Then, we design a shallow hourglass network [60] to extract the high-resolution feature maps $F_{h,i}$ that preserve the resolution of human appearance while capturing rich low-level features.
+4.1.1. Human-focused Quality Feature Extraction
+For frame ${\bm{y}}_{i}$, we first apply a human body-part segmentation model (Khirodkar et al., 2024), to explicitly detect the body-part masks $M_{i}$, where $M_{i}\in\mathbb{R}^{C_{m}\times H\times W}$ represents six segmentation masks corresponding to the face, arms, torso, legs, feet.
+Then, we design a shallow hourglass network (Newell et al., 2016) to extract the high-resolution feature maps $F_{h,i}$ that preserve the resolution of human appearance while capturing rich low-level features.
 Here, $F_{h,i}\in\mathbb{R}^{C_{h}\times H\times W}$, with $C_{h}$ representing the number of channels of $F_{h,i}$.
 To focus on features specific to human body regions, we multiply the feature maps $F_{h,i}$ with the body masks $M_{i}$ along the channel dimensions to derive the human body-aware feature maps $F_{{\rm body},i}$:
-$F_{{\rm body},i}=F_{h,i}\cdot M_{i},$ (1)
+(1) $F_{{\rm body},i}=F_{h,i}\cdot M_{i},$
 where $F_{{\rm body},i}\in\mathbb{R}^{C_{m}\times C_{b}\times H\times W}$.
@@
 To achieve this, we aggregate the feature maps $F_{{\rm body},i}$ across the channel dimension $C_{h}$:
-$F_{{\rm inner},i}=\sum_{j=0}^{C_{h}-1}F_{{\rm body},i}[,j:,:,:],$ (2)
+(2) $F_{{\rm inner},i}=\sum_{j=0}^{C_{h}-1}F_{{\rm body},i}[,j:,:,:],$
 where $F_{{\rm inner},i}\in\mathbb{R}^{C_{m}\times H\times W}$ contains $C_{m}$ feature maps, each focusing on quality analysis for one of the $C_{m}$ body parts.
@@
 To achieve this, we aggregate the feature maps $F_{{\rm body},i}$ across the channel dimension $C_{m}$:
-$F_{{\rm cross},i}=\sum_{j=0}^{C_{m}-1}F_{{\rm body},i}[j,:,:,:],$ (3)
+(3) $F_{{\rm cross},i}=\sum_{j=0}^{C_{m}-1}F_{{\rm body},i}[j,:,:,:],$
 where $F_{{\rm cross},i}\in\mathbb{R}^{C_{s}\times H\times W}$ contains $C_{s}$ feature maps focused on the quality analysis of all body parts.
-Feature Refinement and Fusion. For $F_{{\rm inner},i}$ and $F_{{\rm cross},i}$, we apply a CNN followed by a Transformer encoder [23] to further refine the inner-body and cross-body features:
-$F^{\prime}_{{\rm inner},i}={\rm Transformer}({\rm Conv}(F_{{\rm inner},i})),$ (4)
+Feature Refinement and Fusion.
+For $F_{{\rm inner},i}$ and $F_{{\rm cross},i}$, we apply a CNN followed by a Transformer encoder (Han et al., 2021) to further refine the inner-body and cross-body features:
+(4) $F^{\prime}_{{\rm inner},i}={\rm Transformer}({\rm Conv}(F_{{\rm inner},i})),$
 $F^{\prime}_{{\rm cross},i}={\rm Transformer}({\rm Conv}(F_{{\rm cross},i})),$
 where $\rm Transformer$ denotes the Transformer encoder, and the $\rm Conv$ represents a two-layer CNN.
-Finally, we concatenate them and then apply a CNN layer followed by a self-attention layer [66] to derive the human-focused quality features:
-$F_{{\rm bq},i}$ $={\rm Attention}({\rm Conv}([F^{\prime}_{{\rm inner},i},F^{\prime}_{{\rm cross},i}])),$ (5)
+Finally, we concatenate them and then apply a CNN layer followed by a self-attention layer (Shaw et al., 2018) to derive the human-focused quality features:
+(5) $F_{{\rm bq},i}$ $={\rm Attention}({\rm Conv}([F^{\prime}_{{\rm inner},i},F^{\prime}_{{\rm cross},i}])),$
 where $\rm Attention$ denotes the self-attention operator and $F_{{\rm bq},i}$ is the human-focused quality features.
-4.1.2 Holistic Quality Feature Extraction
-To extract quality-aware features of the entire frame, we pre-train a ViT [15] model on Pick-a-Pic [37], a large-scale text-to-image (T2I) quality assessment dataset, to learn the quality-aware feature representation for AIGC images.
-Then, for frame ${\bm{y}}_{i}$, we employ the pre-trained ViT [15] as the feature extractor to compute the holistic quality features:
-$F_{{\rm hq},i}={\rm{ViT}}({\bm{y}}_{i}).$ (6)
-Table 1: The performance comparison of existing IQA-based, AQA-based, and VQA-based metrics with our proposed GHVQ metrics on the Human-AGVQA dataset. NSC refers to natural scene content.
+4.1.2. Holistic Quality Feature Extraction
+To extract quality-aware features of the entire frame, we pre-train a ViT (Dosovitskiy et al., 2021) model on Pick-a-Pic (Kirstain et al., 2023), a large-scale text-to-image (T2I) quality assessment dataset, to learn the quality-aware feature representation for AIGC images.
+Then, for frame ${\bm{y}}_{i}$, we employ the pre-trained ViT (Dosovitskiy et al., 2021) as the feature extractor to compute the holistic quality features:
+(6) $F_{{\rm hq},i}={\rm{ViT}}({\bm{y}}_{i}).$
+Table 1. The performance comparison of existing IQA-based, AQA-based, and VQA-based metrics with our proposed GHVQ metrics on the Human-AGVQA dataset. NSC refers to natural scene content.
 Quality Dimension Methods Pre-training / Initialization Model Type Zero-shot Fine-tuning
 SRCC PLCC SRCC PLCC
-Human Appearance NIQE (ISPL, 2012) [58] NA (handcraft) NSC 0.234 0.273 0.564 0.583
-BRISQUE (TIP, 2012) [57] NA (handcraft) NSC 0.261 0.314 0.593 0.629
-HyperIQA (CVPR, 2020) [68] TID2013 [63] NSC 0.314 0.339 0.684 0.705
-UNIQUE (TIP, 2021) [101] KonIQ-10K [32] NSC 0.243 0.280 0.671 0.695
-MUSIQ (ICCV, 2021) [34] KonIQ-10K [32] NSC 0.229 0.262 0.584 0.609
-StairIQA (JSTSP, 2023) [72] KonIQ-10K [32] NSC 0.334 0.393 0.665 0.671
-CLIP-IQA (AAAI, 2023) [81] KonIQ-10K [32] NSC 0.329 0.368 0.672 0.683
-LIQE (CVPR, 2023) [102] KonIQ-10k [32] NSC 0.285 0.332 0.641 0.664
-MA-AGIQA (ACMMM, 2024) [84] AGIQA-3k [43] AIGC 0.377 0.380 0.728 0.735
-Q-Align (ICML, 2024) [89] fused  [32][17][96][21] NSC 0.364 0.419
-GHVQ (proposed) AIGC —- 0.812 0.821
-Action Continuity ACTION-NET (ACMMM, 2020) [99] Kinetics [6] NSC 0.197 0.221 0.541 0.551
-USDL (CVPR, 2020) [74] Kinetics [6] NSC 0.207 0.260 0.583 0.582
-CoRe (ICCV, 2021) [97] Kinetics [6] NSC 0.179 0.213 0.560 0.561
-TSA (CVPR, 2022) [92] Kinetics [6] NSC 0.202 0.258 0.607 0.611
-Motion Smoothness (CVPR, 2024) [33] AMT [48] AIGC 0.249 0.30
-Temporal Flickering (CVPR, 2024) [33] RAFT [75] AIGC 0.109 0.238
-Action-Score (CVPR, 2024) [54] VideoMAE V2 [82] AIGC 0.208 0.284
-Flow-Score (CVPR, 2024) [54] RAFT [75] AIGC 0.153 0.255
-GHVQ (proposed) AIGC 0.788 0.792
-Overall Video TLVQM (TIP, 2019) [38] NA (handcraft) NSC 0.275 0.314 0.613 0.624
-RAPIQUE (JSP, 2021) [78] NA (handcraft) NSC 0.313 0.351 0.632 0.648
-VIDEAL (TIP, 2021) [77] NA (handcraft) NSC 0.338 0.357 0.638 0.651
-PatchVQ (CVPR, 2021) [96] LSVQ [96] NSC 0.378 0.404 0.667 0.703
-SimpleVQA (ACMMM, 2022) [71] LSVQ [96] NSC 0.368 0.380 0.695 0.709
-FastVQA (ECCV, 2023) [87] LSVQ [96] NSC 0.385 0.408 0.708 0.733
-DOVER (ICCV, 2023) [88] DIVIDE [88] NSC 0.315 0.337 0.713 0.736
-T2VQA (ACMMM, 2024) [39] T2VQA [39] AIGC 0.357 0.364 0.751 0.754
-UGVQ (Arxiv, 2024) [103] LGVQ [103] AIGC 0.349 0.358 0.748 0.757
-EvalCrafter (CVPR, 2024) [54] DIVIDE [88] NSC 0.328 0.341
-Q-Align (ICML, 2024) [89] fused  [32][17][96][21] AIGC 0.427 0.485
-GHVQ (proposed) AIGC 0.780 0.788
+Human Appearance NIQE (ISPL, 2012) (Mittal et al., 2012b) NA (handcraft) NSC 0.233 0.273 0.564 0.570
+BRISQUE (TIP, 2012) (Mittal et al., 2012a) NA (handcraft) NSC 0.258 0.314 0.594 0.619
+CNNIQA (CVPR, 2014) (Kang et al., 2014) TID2013 (Ponomarenko et al., 2015) NSC 0.243 0.296 0.620 0.646
+HyperIQA (CVPR, 2020) (Su et al., 2020) TID2013 (Ponomarenko et al., 2015) NSC 0.318 0.339 0.684 0.696
+UNIQUE (TIP, 2021) (Zhang et al., 2021) KonIQ-10K (Hosu et al., 2020) NSC 0.239 0.280 0.668 0.684
+MUSIQ (ICCV, 2021) (Ke et al., 2021) KonIQ-10K (Hosu et al., 2020) NSC 0.224 0.262 0.583 0.601
+StairIQA (JSTSP, 2023) (Sun et al., 2023) KonIQ-10K (Hosu et al., 2020) NSC 0.339 0.393 0.661 0.663
+CLIP-IQA (AAAI, 2023) (Wang et al., 2023a) KonIQ-10K (Hosu et al., 2020) NSC 0.327 0.368 0.672 0.674
+LIQE (CVPR, 2023) (Zhang et al., 2023b) KonIQ-10k (Hosu et al., 2020) NSC 0.283 0.332 0.640 0.656
+MA-AGIQA (ACMMM, 2024) (Wang et al., 2024b) AGIQA-3k (Li et al., 2023a) AIGC 0.373 0.380 0.724 0.733
+Q-Align (ICML, 2024) (Wu et al., 2023c) fused  (Hosu et al., 2020)(Fang et al., 2020)(Ying et al., 2021)(Gu et al., 2018) NSC 0.362 0.419 0.725 0.731
+GHVQ (proposed) AIGC —- 0.805 0.809
+Action Continuity ACTION-NET (ACMMM, 2020) (Zeng et al., 2020) Kinetics (Carreira and Zisserman, 2017a) NSC 0.198 0.223 0.541 0.553
+USDL (CVPR, 2020) (Tang et al., 2020) Kinetics (Carreira and Zisserman, 2017a) NSC 0.208 0.259 0.583 0.591
+CoRe (ICCV, 2021) (Yu et al., 2021) Kinetics (Carreira and Zisserman, 2017a) NSC 0.177 0.210 0.562 0.577
+TSA (CVPR, 2022) (Xu et al., 2022) Kinetics (Carreira and Zisserman, 2017a) NSC 0.204 0.256 0.602 0.613
+Motion Smoothness (CVPR, 2024) (Huang et al., 2024) AMT (Li et al., 2023b) AIGC 0.250 0.275
+Temporal Flickering (CVPR, 2024) (Huang et al., 2024) RAFT (Teed and Deng, 2020) AIGC 0.137 0.239
+Action-Score (CVPR, 2024) (Liu et al., 2024a) VideoMAE V2 (Wang et al., 2023b) AIGC 0.209 0.244
+Flow-Score (CVPR, 2024) (Liu et al., 2024a) RAFT (Teed and Deng, 2020) AIGC 0.254 0.279
+GHVQ (proposed) AIGC 0.771 0.778
+Overall Video TLVQM (TIP, 2019) (Korhonen, 2019) NA (handcraft) NSC 0.272 0.312 0.603 0.617
+RAPIQUE (JSP, 2021) (Tu et al., 2021b) NA (handcraft) NSC 0.313 0.351 0.621 0.637
+VIDEAL (TIP, 2021) (Tu et al., 2021a) NA (handcraft) NSC 0.342 0.353 0.628 0.642
+PatchVQ (CVPR, 2021) (Ying et al., 2021) LSVQ (Ying et al., 2021) NSC 0.379 0.399 0.657 0.693
+SimpleVQA (ACMMM, 2022) (Sun et al., 2022) LSVQ (Ying et al., 2021) NSC 0.364 0.378 0.687 0.698
+BVQA (TCSVT, 2022) (Li et al., 2022b) KoNViD-1k (Hosu et al., 2017) NSC 0.349 0.374 0.672 0.702
+FastVQA (ECCV, 2023) (Wu et al., 2022) LSVQ (Ying et al., 2021) NSC 0.390 0.412 0.698 0.711
+DOVER (ICCV, 2023) (Wu et al., 2023b) DIVIDE (Wu et al., 2023b) NSC 0.312 0.337 0.705 0.717
+T2VQA (ACMMM, 2024) (Kou et al., 2024) T2VQA (Kou et al., 2024) AIGC 0.359 0.367 0.737 0.742
+UGVQ (Arxiv, 2024) (Zhang et al., 2024) LGVQ (Zhang et al., 2024) AIGC 0.349 0.358 0.734 0.743
+EvalCrafter (CVPR, 2024) (Liu et al., 2024a) DIVIDE (Wu et al., 2023b) AIGC 0.328 0.336
+Q-Align (ICML, 2024) (Wu et al., 2023c) fused  (Hosu et al., 2020)(Fang et al., 2020)(Ying et al., 2021)(Gu et al., 2018) AIGC 0.422 0.481 0.715 0.723
+GHVQ (proposed) AIGC 0.768 0.773
 Finally, we aggregate the human-focused features and the holistic quality features as the spatial quality features, then average them across all sampled frames:
-$F_{{\rm sq},i}$ $=[F_{{\rm bq},i},F_{{\rm hq},i}],$ (7)
+(7) $F_{{\rm sq},i}$ $=[F_{{\rm bq},i},F_{{\rm hq},i}],$
 $F_{{\rm sq}}$ $=\sum^{Ns-1}_{i=0}F_{{\rm sq},i},$
 where $F_{{\rm sq},i}$ and $F_{{\rm sq}}$ represent the spatial quality features of frame ${\bm{y}}_{i}$ and video sequence ${\bm{y}}$.
-4.2 Action Quality Analyzer
-The action quality analyzer is designed to quantify temporal continuity of videos, serving as a complement to the spatial quality analyzer in assessing comprehensively the quality of AGVs.
+4.2. Action Quality Analyzer
+The action quality analyzer is designed to quantify the temporal continuity of videos, serving as a complement to the spatial quality analyzer in assessing comprehensively the quality of AGVs.
 To address this, we utilize an action recognition model, as the action quality analyzer to capture the temporal continuity of human body movement.
-Since action recognition models are trained on large-scale, authentic-captured human action datasets, such as the Kinetics series [6], they can effectively extract motion representations of human bodies.
-Moreover, previous VQA studies [71] have also demonstrated that these features are useful for video quality evaluation.
-Specifically, for video $\bm{x}$ and the action recognition network $\rm{SlowFast}$ [19], we calculate the action quality features $F_{\rm aq}$ using all frames in $\bm{x}$.
-4.3 Text Feature Extractor
+Since action recognition models are trained on large-scale, authentic-captured human action datasets, such as the Kinetics series (Carreira and Zisserman, 2017a), they can effectively extract motion representations of human bodies.
+Moreover, previous VQA studies (Sun et al., 2022) have also demonstrated that these features are useful for video quality evaluation.
+Specifically, for video $\bm{x}$ and the action recognition network $\rm{SlowFast}$ (Feichtenhofer et al., 2019), we calculate the action quality features $F_{\rm aq}$ using all frames in $\bm{x}$.
+4.3. Text Feature Extractor
 Since the purpose of T2V models is to generate video content that aligns with the text prompt, it is necessary to measure the alignment between the video content and the text prompt.
-Following most T2I and T2V alignment evaluation studies [33, 84], we utilize the text encoder of CLIP [37] to extract semantic features $F_{\rm tq}$ of the text prompt $p$.
-4.4 Quality Regressor
+Following most T2I and T2V alignment evaluation studies (Huang et al., 2024; Wang et al., 2024b), we utilize the text encoder of CLIP (Kirstain et al., 2023) to extract semantic features $F_{\rm tq}$ of the text prompt $p$.
+4.4. Quality Regressor
 The quality regressor is used to assess the three quality scores and determine whether human bodies are present and if there exist semantic distortions.
 So, we first integrate the spatial quality features, the action quality features, and the text features into the final quality-aware features $F_{q}$:
-$F_{q}=[F_{\rm sq},F_{\rm aq},F_{\rm tq}].$ (8)
+(8) $F_{q}=[F_{\rm sq},F_{\rm aq},F_{\rm tq}].$
 Then, $F_{q}$ are fed into two multi-layer perceptron (MLP) to produce the quality scores $\hat{q}$ and distortion labels $\hat{b}$.
-The loss function consists of the mean absolute error (MAE) loss, rank loss [39], and binary cross-entropy loss.
-The MAE loss, and rank loss are used to optimize the quality scoring task, and the binary cross-entropy loss is used to optimize the distortion identification task.
-5 Experiments
-5.1 Experiment Settings
+The loss function consists of the mean absolute error (MAE) loss, rank loss (Wen and Wang, 2021), and binary cross-entropy loss.
+The MAE loss and rank loss are used to optimize the quality scoring task, and the binary cross-entropy loss is used to optimize the distortion identification task.
+5. Experiments
+Table 2. The accuracy of MLMM models in identifying the occurrences and distortions of body parts in the Human-AGVQA dataset.
+Models Occurrences / Distortions
+Face Hands Arms Torso Legs Feet Avg
+LLaMAVID (7B) (Li et al., 2025) 83.25 / 38.24 26.94 / 31.99 68.61 / 36.00 81.72 / 46.40 57.91 / 30.60 24.99 / 21.39 57.24 / 34.10
+VideoChatGPT (7B) (Maaz et al., 2023) 67.11 / 61.58 20.09 / 51.13 51.75 / 49.65 64.52 / 37.54 41.36 / 22.93 29.11 / 19.26 45.66 / 40.35
+VideoLLaMA2 (7B) (Cheng et al., 2024) 26.70 / 24.77 20.64 / 14.82 21.64 / 20.29 54.41 / 31.13 21.74 / 45.72 30.33 / 48.24 29.24 / 30.83
+VILA1.5 (7B) (Lin et al., 2024) 35.65 / 26.52 23.46 / 22.39 29.79 / 27.38 57.82 / 30.63 27.52 / 41.63 31.19 / 46.66 34.24 / 32.54
+NeXT-Video (7B) (Li et al., 2024b) 60.97 / 42.56 35.96 / 27.21 45.24 / 31.18 64.48 / 51.80 32.22 / 34.85 45.23 / 57.53 47.35 / 40.86
+OneVison (7B) (Li et al., 2024a) 62.21 / 44.06 39.21 / 39.05 46.17 / 33.50 66.90 / 53.13 33.27 / 40.93 47.11 / 50.17 49.15 / 43.47
+Qwen2-VL (7B) (Wang et al., 2024a) 72.90 / 52.92 47.88 / 31.59 59.08 / 36.83 71.03 / 55.43 47.12 / 65.97 54.01 / 64.84 58.67 / 51.26
+Qwen2.5-VL (7B) (Bai et al., 2025) 87.66 / 68.38 65.15 / 38.79 72.13 / 47.18 81.81 / 63.31 59.54 / 71.04 69.00 / 72.77 72.55 / 60.25
+DeepSeek-VL2 (4.1B) (Wu et al., 2024) 71.80 / 53.56 44.81 / 33.38 57.45 / 51.60 82.67 / 50.67 43.89 / 62.97 60.95 / 52.76 60.26 / 50.82
+GPT-4o  (OpenAI, 2024a) 92.33 / 29.63 82.28 / 24.95 88.51 / 19.23 84.95 / 60.31 62.04 / 73.31 75.96 / 71.02 81.01 / 46.41
+GPT-4o-mini  (OpenAI, 2024a) 90.50 / 28.29 80.35 / 22.60 86.41 / 20.03 83.55 / 61.49 62.52 / 72.24 73.94 / 69.16 79.55 / 45.64
+GHVQ (proposed) 92.99 / 78.36 84.78 / 62.86 88.50 / 70.25 90.23 / 73.13 64.26 / 83.71 77.37 / 87.35 83.02 / 75.94
+5.1. Experiment Settings
 Compared Quality Metrics.
-Since no specific metrics have been designed to evaluate human appearance quality, we compared $10$ IQA models to assess the quality of video frames as a proxy for human appearance.
-For action continuity quality and overall video quality, we select $8$ action quality assessment (AQA) and $11$ VQA methods as comparison respectively.
-For body parts detection and their distortion identification, there are also no specific metrics, we benchmark $9$ general video large multi-modality (LMM) model for this task.
+Since no specific metrics have been designed to evaluate human appearance quality, we compared $11$ IQA models to assess the quality of video frames as a proxy for human appearance.
+For action continuity quality and overall video quality, we select $8$ action quality assessment (AQA) and $12$ VQA methods as comparison respectively.
+For body parts detection and their distortion identification, there are also no specific metrics, we benchmark $11$ general video multi-modal large multi-modality (MLMM) model for this task.
 A detailed introduction to these methods can be found in the supplemental material.
@@
 For video quality scoring, the compared metrics are evaluated on the test set using two approaches: zero-shot testing and fine-tuning testing.
-For body parts detection and distortion identification, we directly test zero-shot performance of compared LMMs.
+For body parts detection and distortion identification, we directly test zero-shot performance of compared MLMMs.
 Evaluation Criteria.
@@
 Other experimental settings including the training details of GHVQ can be found in the supplemental material.
-5.2 Performance Comparison
-5.2.1 Quality Scoring Task
+5.2. Performance Comparison
+5.2.1. Quality Scoring Task
 Human Appearance Quality.
@@
 This may be because these methods focus on learning feature representations from global images, rather than human-focused regions.
-The proposed GHVQ achieves the best performance, outperforming the second-best method by $0.812$ and $0.821$ in terms of SRCC and PLCC, demonstrating the effectiveness of its design for extracting human appearance-focused features.
-Table 2: The accuracy of video LMM models in identifying the occurrences and distortions of body parts in the Human-AGVQA dataset.
-Models Occurrences / Distortions
-Face Arms Torso Legs Feet Avg
-LLaMAVID (7B) [47] 91.9 / 56.6 82.9 / 62.1 81.8 / 46.5 37.2 / 30.7 25.1 / 41.5 63.8 / 47.7
-VideoChatGPT (7B) [56] 73.9 / 71.7 79.7 / 76.2 75.2 / 37.6 38.8 / 23.0 29.2 / 19.4 59.4 / 45.6
-VideoLLaMA2 (7B) [11] 22.5 / 24.9 14.1 / 14.9 41.2 / 61.2 61.9 / 75.8 70.4 / 78.3 42.2 / 51.0
-VILA1.5 (7B) [51] 22.4 / 36.6 20.1 / 22.5 16.5 / 60.7 60.8 / 74.7 71.3 / 76.8 38.2 / 54.3
-NeXT-Video (7B) [44] 62.5 / 37.6 56.8 / 22.5 78.6 / 60.9 58.2 / 75.0 65.3 / 77.6 64.3 / 54.7
-OneVison (7B) [42] 61.9 / 29.8 71.9 / 19.8 64.6 / 62.2 52.2 / 77.0 67.2 / 80.3 63.6 / 53.8
-Qwen2-VL (7B) [83] 63.1 / 44.3 57.9 / 25.7 71.0 / 60.4 62.6 / 76.1 74.1 / 79.9 70.3 / 57.3
-GPT-4o  [61] 92.4 / 29.7 89.7 / 19.3 85.1 / 61.6 62.1 / 77.4 76.1 / 80.1 80.0 / 53.6
-GPT-4o-mini  [61] 90.6 / 28.4 86.5 / 20.1 82.7 / 60.4 62.6 / 75.3 74.0 / 77.3 78.6 / 52.3
-GHVQ (proposed) 92.6 / 72.3 87.1 / 80.6 87.1 / 63.4 63.5 / 78.1 77.2 / 81.5 81.5 / 75.2
+The proposed GHVQ achieves the best performance, outperforming the second-best method by $0.805$ and $0.809$ in terms of SRCC and PLCC, demonstrating the effectiveness of its design for extracting human appearance-focused features.
 Action Continuity Quality.
@@
 This is because AVA methods are typically designed to quantify the completeness of professional sports performances, while the temporal metrics in AIGC VQA methods rely on handcrafted motion descriptors that do not capture semantic content, such as those involving individuals.
-In the fine-tuning results, GHVQ outperforms the re-trained AVA methods, surpassing the second-best method by $0.788$ and $0.792$ in terms of SRCC and PLCC, demonstrating that the proposed model architecture is better suited to measure action continuity quality.
+In the fine-tuning results, GHVQ outperforms the re-trained AVA methods, surpassing the second-best method by $0.771$ and $0.778$ in terms of SRCC and PLCC, demonstrating that the proposed model architecture is better suited to measure action continuity quality.
 Overall Video Quality.
@@
 Similarly, re-training these methods leads to significant improvements.
-Since GHVQ comprehensively extracts human-focused, action continuity, and holistic frame quality features, the proposed GHVQ outperforms existing VQA methods, improving performance by $0.780$ and $0.788$ in terms of SRCC and PLCC.
-Table 3: Ablation study results for different modules across quality metrics.
+Since GHVQ comprehensively extracts human-focused, action continuity, and holistic frame quality features, the proposed GHVQ outperforms existing VQA methods, improving performance by $0.768$ and $0.773$ in terms of SRCC and PLCC.
+5.2.2. Body Presence and Distortion Identifying Task
+Identifying the presence and distortions of human body parts is a novel task, and we benchmark the performance of MLMM models for this task in Table 2.
+The key prompts that we utilized for the evaluation are listed below.
+Note that replace [body part] with one of the following: ”face”, ”hands”, ”arms”, ”torso”, ”legs”, or ”feet”.
+More details are shown in the supplementary material.
+For body part occurrences:
+# Is there a person’s [body part] appearing in this video? Answer with ”Yes” or ”No”. #.
+For body part distortions:
+# Is there any incompleteness, unrealistic appearance, or discontinuous movements of the person’s [body part] in this video? Answer with ”Yes” or ”No”. #.
+As for MLMMS, the GPT-4 series performs best in terms of detecting the presence of body parts, while Qwen2.5-VL achieves the highest accuracy in identifying whether body parts exhibit semantic distortions.
+We observe that that identifying distortions in arms is the most challenging, followed by facial distortions while detecting distortions in legs and feet is relatively easier.
+This aligns with the inherent complexity of structures like faces and arms (particularly hands).
+Overall, our GHVQ model outperforms in both presence detection and distortion identification tasks, improving accuracy by $2.01\%$ and $15.69\%$, respectively.
+Table 3.
+Ablation study results for different modules across quality metrics.
 SQA, AQA, and TFE refer to the spatial quality analyzer, action quality analyzer, and text feature extractor, respectively.
@@
 SQA AQA TFE HA Quality AC Quality OV Quality Distortion Identification
-0.633 / 0.638 0.617 / 0.631 0.724 / 0.733 0.575 / 0.512
-0.767 / 0.770 0.747 / 0.748 0.728 / 0.740 0.753 / 0.704
-0.781 / 0.789 0.754 / 0.765 0.773 / 0.776 0.802 / 0.736
-0.812 / 0.821 0.788 / 0.792 0.780 / 0.788 0.815 / 0.752
-5.2.2 Body Presence and Distortion Identifying Task
-Identifying the presence and distortions of human body parts is a novel task, and we benchmark the performance of video LMM models for this task in Table 2.
-As for LMMS, GPT-4 series perform best in terms of detecting the presence of body parts, while Qwen2-VL achieve the highest accuracy in identifying whether body parts exhibit semantic distortions.
-We observe that that identifying distortions in arms is the most challenging, followed by facial distortions while detecting distortions in legs and feet is relatively easier.
-This aligns with the inherent complexity of structures like faces and arms (particularly hands).
-Overall, our GHVQ model outperforms in both presence detection and distortion identification tasks, improving accuracy by $1.8\%$ and $28.3\%$, respectively.
-5.3 Ablation Study
-We conducted an ablation study to validate the effectiveness of the spatial quality analyzer, action quality analyzer, and text feature extractor.
-The experimental results are shown in Table 3.
+0.625 / 0.631 0.611 / 0.626 0.717 / 0.730 58.36 / 52.16
+0.762 / 0.767 0.741 / 0.742 0.721 / 0.734 75.94 / 71.08
+0.774 / 0.782 0.752 / 0.761 0.760 / 0.765 81.02 / 73.91
+0.805 / 0.809 0.771 / 0.778 0.768 / 0.773 83.02 / 75.94
+5.3. Ablation Study
+We conducted an ablation study to validate the effectiveness of the spatial quality analyzer, action quality analyzer, and text feature extractor. The experimental results are shown in Table 3.
 We observe that the performance across all three quality dimensions and distortion identification degrades the most when the spatial quality analyzer is removed, highlighting the importance of extracting human-focused and AIGC-aware quality features.
@@
 In summary, the combination of all three modules achieves the highest performance across all quality dimensions.
-6 Conclusion
+6. Conclusion
 In this paper, we present an in-depth quality assessment study on human activity AGVs, which includes the construction of the large-scale human activity AGV dataset Human-AGVQA and the development of the objective quality metric GHVQ for human activity AGVs.
@@
 We hope these contributions will promote the development and application of T2V models.
-Supplementary Material
-In this supplementary file, we provide more details of Human-AGVQA dataset in Section 7, including text prompts selection, T2V model details, and subjective assessment experiments.
-We then provide data processing details and further analysis in Section 8, including MOS analysis for T2V models, distortion identification analysis, and the relationship between distortion identification and visual quality.
-The detailed experiment results are discussed in Section 9, including evaluation criteria in our experiments, compared quality metrics in the proposed benchmark, training details of proposed GHVQ metric, and ablation study of different backbones.
-7 Human-AGVQA Dataset
-7.1 Text Prompts Selection
-Figure 7: The distribution of 44 subcategories in 8 categories of the 400 prompts.
+Appendix A Human-AGVQA Dataset
+A.1. Text Prompts Selection
+Figure 7. The distribution of 44 subcategories in 8 categories of the 400 prompts.
 To ensure that the text prompts in our dataset encompass a wide range of real-world human activities, we perform a comprehensive and systematic classification of the words used to create human activity-oriented text prompts.
-Specifically, we first divide the words into $8$ categories according to their properties: ages, genders, races, occupations, scenes, emotion, appearance, and activities.
+Specifically, the categories of human activities are derived from the AIGC VQA dataset, Vbench(Huang et al., 2024), FETV(Liu et al., 2023), and real-world video caption datasets, InternVid-10M(Radford et al., 2021a) and Webvid-10M(Bain et al., 2021).
+We use GPT-4 to identify high-frequency activities within these datasets, and finally we divide the words into $8$ categories according to their properties: ages, genders, races, occupations, scenes, emotion, appearance, and activities.
 Each of these categories is then further subdivided into more specific subcategories that are commonly observed in daily human life, as illustrated in Table 4.
-Some example text prompts and their corresponding generated videos are illustrated in Figure 9.
+The prompts require at least one person to be present.
+60% of the prompts specify a single person, while 40% involve multiple individuals.
 The proportions of each subcategory within their respective categories are illustrated in Figure 7.
@@
 The average text prompt length is $13.85$, with a median of $14$ words.
-Figure 8: Histogram and density plot of word count per prompt.
-7.2 T2V models
-Detailed information about eight T2V methods is presented in Table 5.
-All selected T2V models have been published within the past two years and represent the most advanced video generation capabilities of T2V models to date.
-Except for VideoCrafter2, the resolution of videos generated by these models is at least $512$, indicating high clarity.
-However, the frame rates of StreamingT2V, Latte, and AnimateDiff are below $10$ fps, which may result in low motion consistency in actions.
-7.3 Subjective Quality Assessment Experiments
-7.3.1 Visual Quality Scoring Criteria
-Table 4: Subcategories of ages, genders, races, occupations, scenes, emotion, appearance, and activities.
+Table 4. Subcategories of ages, genders, races, occupations, scenes, emotion, appearance, and activities.
 Category Subcategories
@@
 activity study, work, socializing, entertainment, wilderness, sports
+Figure 8. Histogram and density plot of word count per prompt.
+Table 5. Video formats generated by the $15$ T2V models in the Human-AGVQA dataset.
+Methods Duration (s) FPS Resolution Launch Month Utilized Version Version Release Month Open Source
+Show-1 (Zhang et al., 2023a) $3.4$ $8$ $1914\times 951$ 2023.10 - 2023.10
+VideoCrafter2 (Chen et al., 2024a) $1.6$ $10$ $512\times 320$ 2023.04 - 2024.01
+StreamingT2V (Henschel et al., 2024) $5.0$ $8$ $1280\times 720$ 2024.03 - 2024.03
+AnimateDiff (Guo et al., 2024) $2.0$ $8$ $512\times 512$ 2024.03 AnimateDiff-Lightning 2024.03
+MagicTime (Yuan et al., 2024) $2.1$ $23$ $512\times 512$ 2024.04 - 2024.04
+Open-sora-plan (Lab and etc., 2024) $2.7$ $24$ $512\times 512$ 2024.03 v1.1.0 2024.05
+Latte (Ma et al., 2024) $2.0$ $8$ $512\times 512$ 2024.01 Latte-1 2024.05
+Gen-2 (Esser et al., 2023) $4.0$ $24$ $1408\times 768$ 2023.02 Gen-2 2024.06
+StableVideo (AI, 2024a) $4.0$ $24$ $1024\times 576$ 2023.11 - 2024.06
+Dreamina (ByteDance, 2024) $5.0$ $24$ $1280\times 951$ 2024.05 S2.0 2024.11
+LTX-Video (HaCohen et al., 2024) $4.8$ $25$ $704\times 480$ 2024.11 v0.9.1 2024.12
+Ying (AI, 2024b) $5.0$ $30$ $960\times 720$ 2024.07 - 2024.12
+Sora (OpenAI, 2024b) $5.0$ $24$ $512\times 512$ 2024.02 480p 2024.12
+Kling (Technology, 2024) $4.0$ $24$ $1024\times 576$ 2024.06 kling1.6 2024.12
+Wan (Cloud, 2025) $5.3$ $30$ $1280\times 720$ 2023.07 wanx2.1-t2v-turbo 2025.01
+A.2. T2V models
+Table 5 provides detailed information about 15 T2V models, including 7 commercially available ones.
+These models are sorted by the release time of the version used.
+All selected models were released within the past two years, showcasing the state-of-the-art capabilities in video generation for T2V models.
+The video resolution of most models is at least 512x512, ensuring high clarity.
+Notably, VideoCrafter2 and LTX-Video generate videos with lower resolutions.
+In terms of frame rate, some models such as Show-1, StreamingT2V, Latte, and AnimateDiff produce videos with frame rates below 10 fps.
+This could lead to potential issues with motion consistency in actions, as lower frame rates may cause choppy or less fluid motion.
+Some example text prompts and their corresponding generated videos are illustrated in Figure 9.
+A.3. Subjective Quality Assessment Experiments
+A.3.1. Visual Quality Scoring Criteria
 The quality scores of AGVs are rated from three dimensions: human appearance quality, action continuity quality, and overall video quality.
@@
 If the video does not contain any human body parts (despite the prompt requiring them), this score is set to 1.
-Table 5: Video formats generated by the eight T2V models in the Human-AGVQA dataset.
-Methods Open Source Year Duration (s) FPS Resolution
-Latte [55] 2023 $2.0$ $8$ $512\times 512$
-AnimateDiff [22] 2023 $2.0$ $8$ $512\times 512$
-Gen-2 [16] 2023 $4.0$ $24$ $1408\times 768$
-VideoCrafter2 [9] 2024 $1.6$ $10$ $512\times 320$
-MagicTime [98] 2024 $2.1$ $23$ $512\times 512$
-StreamingT2V [26] 2024 $5.0$ $8$ $1280\times 720$
-Open-sora-plan [40] 2024 $2.7$ $24$ $512\times 512$
-StableVideo [1] 2024 $4.0$ $24$ $1024\times 576$
-Figure 9: Some example frames of $8$ text categories and $8$ T2V models in proposed Human-AGVQA dataset.
-Figure 10: Annotation Interface for Subjective Quality Experiment.
+Figure 9. Some example frames of $8$ text categories and $8$ T2V models in proposed Human-AGVQA dataset.
 – Action Continuity Quality
@@
 The video is barely watchable and lacks any sense of coherence.
-7.3.2 Semantic Distortion Identification Criteria
+A.3.2. Semantic Distortion Identification Criteria
 – Human Body Presence
@@
 For example, facial features are proportionate, arm movements are smooth and natural, the torso connects properly with other body parts, and the shapes and motions of the legs and feet align with expected behavior.
-Figure 11: Annotation Interface for Distortion Identification.
-7.3.3 Subjective Quality Experiment Procedure
+Figure 10. Annotation Interface for Subjective Quality Experiment.
+Figure 11. Annotation Interface for Distortion Identification.
+A.3.3. Subjective Quality Experiment Procedure
 A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years.
@@
 Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification.
-Each video was rated by $10$ subjects, and labeled by $5$ experts.
+Each video was rated by $12$ subjects, and labeled by $5$ experts.
 The annotation interface for subjective quality experiment and distortion identification are shown in Figure 10 and Figure 11.
 All subjects had normal or corrected-to-normal vision.
-The experiments were conducted in a controlled environment according to the recommendations of ITU-R BT.500-13 [2] to minimize external variables that could influence the judgments of the subjects.
+The experiments were conducted in a controlled environment according to the recommendations of ITU-R BT.500-13 (Met, 2002) to minimize external variables that could influence the judgments of the subjects.
 The setup included 27-inch calibrated display monitors with 95% DCI-P3 color gamut and a resolution of 4K.
@@
 This session aimed to familiarize them with the evaluation criteria and the rating interface.
-In the formal experiment, $3,200$ videos were divided into eight groups, each containing $400$ videos that covered all 400 prompts.
+In the formal experiment, $6,000$ videos were divided into $15$ groups, each containing $400$ videos that covered all $400$ prompts.
 To avoid visual fatigue, each session lasted no longer than $30$ minutes, ensuring participants could maintain a high level of attention and accuracy in their ratings.
-In total, there are $96,000$ opinion scores and $160,000$ binary labels in the Human-AGVQA dataset.
-8 Data Processing and Analysis
-We follow the recommended method in  [2] to process the subjective ratings collected during the experiment.
+In total, there are $270,000$ opinion scores and $360,000$ binary labels in the Human-AGVQA dataset.
+Appendix B Data Processing and Analysis
+We follow the recommended method in  (Met, 2002) to process the subjective ratings collected during the experiment.
 Outlier ratings are detected and removed if they deviate by more than $2\sigma$ (for normal distributions) or $\sqrt{20}\sigma$ (for non-normal distributions) from the mean rating for that condition.
@@
 For each test condition, the mean score ($\mu_{i}$) and standard deviation ($\sigma_{i}$) are calculated based on all valid ratings provided by observer $i$:
-$\mu_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}s_{ij},\quad\sigma_{i}=\sqrt{\frac{1}{N_{i}-1}\sum_{j=1}^{N_{i}}(s_{ij}-\mu_{i})^{2}}$ (9)
+(9) $\mu_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}s_{ij},\quad\sigma_{i}=\sqrt{\frac{1}{N_{i}-1}\sum_{j=1}^{N_{i}}(s_{ij}-\mu_{i})^{2}}$
 where $s_{ij}$ represents the raw rating assigned by observer $i$ to condition $j$, and $N_{i}$ is the number of conditions rated by observer $i$.
 To mitigate individual bias, each raw score $s_{ij}$ is normalized to a Z-score:
-$Z_{ij}=\frac{s_{ij}-\mu_{i}}{\sigma_{i}}$ (10)
+(10) $Z_{ij}=\frac{s_{ij}-\mu_{i}}{\sigma_{i}}$
 Finally, the Mean Opinion Score (MOS) for each test condition $j$ is computed as the average of the normalized Z-scores across all observers ($M_{j}$):
-$MOS_{j}=\frac{1}{M_{j}}\sum_{i=1}^{M_{j}}Z_{ij}$ (11)
+(11) $MOS_{j}=\frac{1}{M_{j}}\sum_{i=1}^{M_{j}}Z_{ij}$
 For distortion identification, we use a voting method to determine body presence and body distortion. The label with the highest number of votes is selected as the final result.
-8.1 Inter-subject Consistency
-To evaluate the consistency and reliability of quality scoring, we compute the inter-annotator agreement metric, Krippendorff’s Alpha ($\alpha$), where Krippendorff’s Alpha ($\alpha$) for HA quality, AC quality, OV quality are $0.708$, $0.695$, and $0.628$, respectively, indicating appropriate variations among subjects.
-For semantic distortion identification, we calculated the proportions of options receiving 3 votes, 4 votes, and 5 votes, which were $9.7\%$, $41.5\%$, and $46.4\%$, respectively.
+B.1. Inter-subject Consistency
+To evaluate the consistency and reliability of quality scoring, we compute the inter-annotator agreement metric, Krippendorff’s Alpha ($\alpha$), where Krippendorff’s Alpha ($\alpha$) for HA quality, AC quality, OV quality are $0.643$, $0.658$, and $0.705$, respectively, indicating appropriate variations among subjects.
+For semantic distortion identification, we calculated the proportions of options receiving 3 votes, 4 votes, and 5 votes, which were $9.1\%$, $44.6\%$, and $43.8\%$, respectively.
 This demonstrates that the majority of the votes from subjects were consistent.
-8.2 MOS Analysis for Each T2V Model
-We analyzed the MOS of three quality dimensions across $44$ attributes (mentioned in Table 4) for eight T2V models (mentioned in Table 5), as shown in Figure 12.
-Figure 12: The MOS comparison for $8$ text categories of each T2V model.
+B.2. MOS Analysis for Each T2V Model
+We analyzed the MOS of three quality dimensions across $44$ attributes (mentioned in Table 4) for $15$ T2V models (mentioned in Table 5), as shown in Figure 12.
+Figure 12. The MOS comparison for $8$ text categories of each T2V model.
+Sora.
+Sora excels in all three aspects.
+Overall video (OV) quality emerges as the strongest aspect of Sora, consistently achieving high scores on most attributes, including complex scenarios such as workplace in scene and emotions like anger.
+This highlights the robust capacity of the model to generate visually coherent output.
+Human appearance (HA) quality also performs well.
+In contrast, action continuity (AC) quality represents the model’s primary weakness in most cases.
+This is particularly evident in subcategories requiring complex movements or high dynamism, such as anger in emotion and sports in activity, revealing the model’s limitations in modeling motion continuity.
+Show1.
+Show1 demonstrates average performance across the three evaluated dimensions.
+Notably, the AC Quality is slightly higher than the HA Quality in most cases.
+Additionally, Show1 maintains a relatively balanced performance across various attribute prompts.
+This indicates that the model is capable of handling a wide range of prompts with consistent results, without significant bias toward any particular category.
+However, there are slight performance drops in specific areas, such as the athlete in occupation, disgust in emotion, and sports in activity.
+Dreamina.
+Dreamina performs well across all three evaluation dimensions.
+Overall, the OV Quality is slightly higher than the other two quality metrics, indicating that the model may face challenges in generating consistent representations of human features and actions.
+Additionally, Dreamina shows noticeable performance discrepancies across different attribute prompts.
+For example, there is a clear distinction between the chef and cleaner in occupations, and the model’s representation of fear in emotions and sports in activities is relatively weaker compared to other categories.
+These variations suggest that it struggles with consistently representing certain categories.
+Ying.
+Ying demonstrates above-average performance across all three evaluation dimensions.
+In most cases, the AC quality slightly exceeds the other two quality metrics.
+However, Ying shows noticeable variability in performance across different attribute prompts, particularly in videos involving significant movement.
+For example, the model’s representation of the athlete in occupation and sports in activity is notably weaker compared to other categories.
+In contrast, the performance across other attributes shows minimal variation.
+This suggests that while the model performs well overall, it faces challenges in consistently representing dynamic actions and high-motion scenarios.
+LTX LTX shows below-average performance across all three evaluation dimensions.
+The HA quality is generally lower than the other metrics, suggesting difficulty in accurately depicting human features.
+Performance varies significantly across attributes, with relatively better results for the middle-aged in age, while other age categories and the painter and student in occupations show poor results.
+The model also struggles with representing excitement and fear in emotions, as well as sports and study in activities.
+These discrepancies highlight the model’s challenges in providing consistent and accurate representations across various attributes.
+Wan Wan performs exceptionally well across all three evaluation dimensions.
+In most cases, the OV quality is slightly higher than the other two metrics.
+Additionally, Wan demonstrates a relatively balanced performance across various attribute prompts, with only slight weaknesses in the athlete in occupation and sports in activity.
+Performance in other attributes shows no significant differences.
+This indicates that Wan generally maintains consistent quality, with room for improvement in dynamic or high-motion scenarios.
+Kling Kling performs exceptionally well across all three evaluation dimensions.
+In most cases, similar to Wan, the OV quality is slightly higher than the other two metrics.
+Kling exhibits a more balanced performance across a wide range of attribute prompts, with only slight weaknesses in the athlete in occupation and sports in activity.
+Performance in other attributes shows no significant variations, demonstrating strong and consistent quality across all areas.
+This indicates that Kling delivers reliable results across different prompts.
 StableVideo.
-Overall video (OV) quality emerges as the strongest aspect of StableVideo, consistently achieving high scores on most attributes, including complex scenarios such as workplace and emotion like anger.
-This highlights the model’s robust capability to generate visually coherent outputs.
-Human appearance (HA) quality also performs well, though it slightly lags behind overall video quality, indicating the model’s ability to generate clear and recognizable human features but with room for improvement compared to its overall visual performance.
-In contrast, action continuity (AC) quality represents the model’s primary weakness, consistently scoring the lowest among the three dimensions.
-This is particularly evident in subcategories requiring complex movements or high dynamism, such as surprise in emotion and sports in activity, revealing the model’s limitations in modeling motion continuity.
+StableVideo performs above average across all three evaluation dimensions.
+However, its HA and AC qualities are noticeably lower than OV quality, suggesting challenges in accurately representing human features and maintaining smooth, continuous actions.
+StableVideo also shows significant variability across different attribute prompts.
+It performs poorly in the venue in scene, the cleaner in occupation, the disgust in emotion, and sports activity.
+These discrepancies indicate that the model struggles with generating complex scenes, effectively conveying emotions, and accurately representing dynamic actions.
+Open-Sora.
+The performance of OV Quality and AC Quality of Open-Sora is not satisfactory, and they still fall short significantly compared to other models.
+In addition, HA quality consistently lags behind, particularly in the occupation, emotion, and activity categories, where the model struggles to render detailed and realistic human appearances.
+These results suggest that it faces challenges in human appearance generation.
 AnimateDiff.
-The AnimateDiff model excels in AC Quality, particularly in emotion and activity, where it consistently outperforms the other quality dimensions.
-This demonstrates the model’s strength in generating continuous and coherent motion.
-OV Quality shows stable and relatively high performance across most attributes, reflecting the model’s ability to maintain overall visual consistency.
-However, HA Quality falls behind, especially in categories such as occupation and activity, where the model struggles with rendering detailed and realistic human appearances.
-These findings suggest that while AnimateDiff is highly competent in motion continuity, it faces challenges in producing visually intricate details.
+AnimateDiff demonstrates moderate performance across all three evaluation dimensions.
+In most cases, its HA quality is significantly lower than its AC quality.
+AnimateDiff’s performance in the athlete in occupation and sports in activity is particularly poor, showing clear discrepancies compared to other categories.
+These findings suggest that while the model performs adequately in certain areas, it struggles with dynamic actions and complex occupations.
+VideoCrafter2.
+VideoCrafter2 shows balanced performance in OV and AC quality, with stable scores across most attributes.
+However, HA quality consistently underperforms, particularly in emotion such as anger and disgust, and activity such as sports and work, where it significantly trails OV and AC quality.
+These results highlight the model’s limitations in generating detailed and realistic human appearances.
+MagicTime.
+MagicTime performs poorly in all three qualities, and there is a certain gap compared with other models.
+In addition, HA quality consistently underperforms, especially in categories involving complex emotions, movements, or appearances, highlighting the model’s limitations in rendering detailed human features.
+The model exhibits significant variability in occupation and appearance, indicating its inconsistent generation quality depending on the attribute.
+Latte.
+Latte shows below-average performance in all three qualities, but it is still slightly insufficient compared with other models.
+OV Quality is the weakest dimension, particularly in occupation and activity categories, where the model demonstrates limitations in generating detailed and realistic human appearances.
+These observations suggest that the Latte model performs moderately well in maintaining overall video quality, and its capacity to generate fine-grained human details requires further improvement.
 Gen-2.
-The Gen-2 model demonstrates exceptional performance in OV Quality, consistently achieving the highest scores across all subcategories.
-This underscores the model’s strength in generating visually cohesive outputs.
-HA Quality ranks second, showing a stable and strong ability to generate clear and recognizable human features, although it slightly trails OV Quality.
-AC Quality, however, is the weakest of the three dimensions, particularly in subcategories involving complex movements, such as angry in emotion and sports in activity, where it exhibits noticeable limitations.
-Latte.
-Latte shows only average performance in terms of OV quality, but it is still slightly insufficient compared with other models.
-AC Quality exhibits greater variability, with strong performance in categories like sadness and surprise, but struggles in more complex occupational categories.
-HA Quality is the weakest dimension, particularly in occupation and activity categories, where the model demonstrates limitations in generating detailed and realistic human appearances.
-These observations suggest that the Latte model performs moderately well in maintaining overall video quality, its capacity to generate fine-grained human details requires further improvement.
-MagicTime.
-MagicTime performs poorly in AC quality among the three quality scores, and there is a certain gap compared with other models.
-In addition, HA quality consistently underperforms, especially in categories involving complex emotions, movements, or appearances, highlighting the model’s limitations in rendering detailed human features.
-The model shows balanced performance in categories like scene and gender but exhibits significant variability in occupation and appearance, indicating its inconsistent generation quality depending on the attribute.
-Open-Sora.
-The performance of OV Quality and AC Quality is not satisfactory, and they still fall short significantly compared to other models.
-In addition, HA quality consistently lags behind, particularly in the occupation, emotion, and activity categories, where the model struggles to render detailed and realistic human appearances.
-These results suggest that while Open-Sora excels in visual consistency and motion continuity, it faces challenges in human appearance generation.
+Gen-2 performs above average across all three evaluation dimensions.
+However, its AC quality is generally lower than the other metrics, indicating challenges in maintaining smooth actions.
+The model also shows significant performance variability, particularly in the athlete and cleaner in occupations, as well as the sports in activity, where its performance is notably weaker.
+This suggests difficulties in accurately representing dynamic actions and certain occupations.
 StreamingT2V.
-StreamingT2V shows only average performance in terms of OV quality and HA quality, indicating its limited capability to generate sufficiently detailed human appearances.
-However, it falls short in AC quality, particularly in dynamic categories such as occupation (Athlete and Cleaner and emotion (surprise and anger), where the scores are significantly lower.
-These findings indicate that while StreamingT2V is effective in static scenarios, its handling of motion continuity in dynamic contexts requires further improvement.
-VideoCrafter2.
-VideoCrafter2 shows balanced performance in OV quality and AC quality, with stable scores across most attributes.
-However, HA quality consistently underperforms, particularly in emotion categories such as anger and disgust, and activity categories such as sports and work, where it significantly trails OV and AC quality.
-These results highlight the model’s limitations in generating detailed and realistic human appearances.
-Nevertheless, the model demonstrates balanced performance in categories like age, scene, ethnicity, and gender, reflecting its comprehensive generation capability in these areas.
-However, its inconsistent performance in emotion, particularly in attributes like anger and disgust, underscores the challenges it faces in modeling complex emotional details.
-8.3 Distortion Identification Analysis
-Table 6: The total and proportion of occurrences and distortions for body parts in the $400$ videos generated by each model. Red represents the minimum value in a column.
+StreamingT2V shows below-average performance in all three qualities, indicating its limited capability to generate sufficiently detailed human appearances.
+It falls short in AC quality, particularly in dynamic categories such as occupation (athlete and cleaner and emotion (surprise and anger), where the scores are significantly lower.
+These findings indicate that its handling of motion continuity in dynamic contexts requires further improvement.
+B.3. Distortion Identification Analysis
+Table 6.
+The total and proportion of occurrences and distortions for body parts in the $400$ videos generated by each model.
+Red, blue, and green represent the first, the second, and the third minimum values in the column, respectively.
 Models Occurrences / Distortions
-Face Arms Torso Legs Feet None / Total
-Latte [55] 358 / 250 324 / 310 329 / 140 89 / 56 48 / 38 21 / 794
-89.5 / 69.8 81.0 / 95.7 82.3 / 42.6 22.3 / 62.9 12.0 / 79.2 5.2 / 69.2
-AnimateDiff [22] 373 / 245 352 / 316 357 / 88 183 / 67 108 / 63 22 / 779
-93.2 / 65.7 88.0 / 89.8 89.2 / 24.6 45.8 / 36.6 27.0 / 58.3 5.5 / 56.7
-Gen-2 [16] 383 / 250 357 / 329 361 / 134 138 / 86 93 / 72 10 / 871
-95.8 / 65.3 89.2 / 92.2 90.2 / 37.1 34.5 / 62.3 23.3 / 77.4 2.5 / 65.4
-StreamingT2V [26] 361 / 304 366 / 342 348 / 199 170 / 127 105 / 92 15 / 1064
-90.2 / 84.2 91.5 / 93.4 87.0 / 57.2 42.5 / 74.7 26.3 / 87.6 3.7 / 78.8
-VideoCrafter2 [9] 382 / 324 350 / 295 365 / 117 98 / 60 64 / 49 11 / 845
-95.5 / 84.8 87.5 / 84.3 91.2 / 32.1 24.5 / 61.2 16.0 / 76.6 2.8 / 67.1
-Open-sora-plan [40] 379 / 333 376 / 371 367 / 214 134 / 84 68 / 60 9 / 1062
-94.8 / 87.9 94.0 / 98.7 91.7 / 58.3 33.5 / 62.7 17.0 / 88.2 2.2 / 80.2
-MagicTime [98] 320 / 292 309 / 302 303 / 177 248 / 177 216 / 187 66 / 1135
-80.0 / 91.2 77.2 / 97.7 75.7 / 58.4 62.0 / 71.4 54.0 / 86.6 16.5/ 81.3
-StableVideo [1] 393 / 289 378 / 323 381 / 127 128 / 60 72 / 47 2 / 846
-98.3 / 73.5 94.5 / 85.4 95.3 / 33.3 32.0 / 46.9 18.0 / 65.3 0.5 / 62.6
+Face Hands Arms Torso Legs Feet Total None
+Latte (Ma et al., 2024) 358 / 284 298 / 295 292 / 166 329 / 174 89 / 65 48 / 41 1414 / 1025 21
+89.5 / 69.8 74.5 / 95.7 73.0 / 95.7 82.3 / 42.6 22.3 / 62.9 12.0 / 79.2 - / 72.5 5.25
+AnimateDiff (Guo et al., 2024) 373 / 272 312 / 299 340 / 116 357 / 124 183 / 82 108 / 67 1673 / 960 21
+93.3 / 65.7 78.0 / 89.8 85.0 / 89.8 89.2 / 24.6 45.8 / 36.6 27.0 / 58.3 - / 57.4 5.25
+Gen-2 (Esser et al., 2023) 383 / 269 328 / 317 362 / 173 361 / 176 138 / 109 93 / 88 1665 / 1132 10
+95.8 / 65.3 82.0 / 92.2 90.5 / 92.2 90.2 / 37.1 34.5 / 62.3 23.3 / 77.4 - / 68.0 2.5
+StreamingT2V (Henschel et al., 2024) 361 / 330 336 / 331 352 / 245 348 / 258 170 / 150 105 / 98 1672 / 1412 17
+90.2 / 84.2 84.0 / 93.4 88.0 / 93.4 87.0 / 57.2 42.5 / 74.7 26.3 / 87.6 - / 84.5 4.25
+VideoCrafter2 (Chen et al., 2024a) 382 / 337 324 / 311 356 / 88 365 / 150 98 / 67 64 / 53 1589 / 1006 10
+95.5 / 84.8 81.0 / 84.3 89.0 / 84.3 91.2 / 32.1 24.5 / 61.2 16.0 / 76.6 - / 63.3 2.5
+Open-sora-plan (Lab and etc., 2024) 379 / 337 342 / 342 357 / 228 367 / 224 134 / 85 68 / 61 1647 / 1277 9
+94.8 / 87.9 85.5 / 98.7 89.3 / 98.7 91.7 / 58.3 33.5 / 62.7 17.0 / 88.2 - / 77.6 2.25
+MagicTime (Yuan et al., 2024) 320 / 296 267 / 266 299 / 184 303 / 182 248 / 183 216 / 189 1653 / 1300 67
+80.0 / 91.2 66.8 / 97.7 74.8 / 97.7 75.7 / 58.4 62.0 / 71.4 54.0 / 86.6 - / 78.7 16.8
+StableVideo (AI, 2024a) 393 / 297 369 / 334 384 / 109 381 / 154 128 / 69 72 / 52 1727 / 1015 2
+98.3 / 73.5 92.3 / 85.4 96.0 / 85.4 95.3 / 33.3 32.0 / 46.9 18.0 / 65.3 - / 58.8 0.5
+Sora (OpenAI, 2024b) 400 / 3 391 / 85 394 / 25 396 / 9 260 / 8 141 / 5 1982 / 135 0
+100.0 / 0.75 97.7 / 21.7 98.5 / 6.35 99.0 / 2.27 65.0 / 3.08 35.2 / 3.55 - / 6.81 0.0
+Show1 (Zhang et al., 2023a) 382 / 294 355 / 326 372 / 168 380 / 35 128 / 42 70 / 26 1687 / 891 12
+95.5 / 77.0 88.8 / 91.3 93.0 / 45.2 95.0 / 9.21 32.0 / 32.8 17.5 / 37.1 - / 52.8 3.0
+Dreamina (ByteDance, 2024) 390 / 17 387 / 202 387 / 59 387 / 14 251 / 22 150 / 18 1952 / 332 3
+97.5 / 4.36 96.8 / 52.2 96.8 / 15.3 96.8 / 3.62 62.8 / 8.76 37.5 / 12.0 - / 17.0 0.75
+Ying (AI, 2024b) 399 / 32 368 / 225 389 / 48 395 / 14 180 / 15 88 / 12 1819 / 346 1
+99.8 / 8.02 92.0 / 61.1 97.3 / 12.3 98.8 / 3.54 45.0 / 8.33 22.0 / 13.6 - / 19.0 0.25
+LTX (HaCohen et al., 2024) 304 / 196 234 / 226 235 / 162 260 / 67 113 / 59 52 / 35 1198 / 745 89
+76.0 / 64.5 58.5 / 96.6 58.8 / 68.9 65.0 / 25.8 28.3 / 52.2 13.0 / 67.3 - / 62.2 22.3
+Wan (Cloud, 2025) 394 / 3 379 / 111 387 / 18 389 / 3 154 / 3 53 / 1 1756 / 139 1
+98.5 / 0.76 94.8 / 29.3 96.8 / 4.65 97.3 / 0.77 38.5 / 1.95 13.3 / 1.89 - / 7.92 0.25
+Kling (Technology, 2024) 393 / 4 367 / 104 385 / 24 393 / 5 222 / 5 123 / 3 1883 / 145 0
+98.3 / 1.02 91.8 / 28.3 96.3 / 6.23 98.3 / 1.27 55.5 / 2.25 30.8 / 2.44 - / 7.70 0.0
 We calculate the proportion of occurrences and distortions for each body part of AGVs for different T2V models, as detailed in Table 6.
-For body occurrences, StableVideo performs the best, with the highest proportion of main body parts (face, arms, torso) present among these T2V models, and only $0.5\%$ of videos generated by StableVideo lack any human depiction.
-In contrast, MagicTime performs the worst, with $16.5\%$ of videos missing any human body parts.
-In terms of distortions, AnimateDiff and StableVideo exhibit better performance, with VideoCrafter2 and Gen-2 also showing comparable results.
-In contrast, Open-Sora and MagicTime show significant deficiencies, with distortions in certain body parts accounting for over $95$%.
-These findings closely align with the subjective quality scores of corresponding T2V models.
-Moreover, an observation of significant concern is that $16.5\%$ of videos generated by MagicTime lack any depiction of human body parts, indicating a shortfall in the model’s ability to maintain consistency between textual descriptions and visual content.
-StableVideo demonstrates the most robust performance in this regard.
-8.4 The Relationship between Semantic Distortion Identification and Visual Quality Scoring
-Figure 13: The correlation between the semantic distortion identification and subjective quality.
+For body part occurrences, Sora and Kling perform exceptionally well, with both models including at least one body part in every generated video.
+Wan, Ying, StableVideo, and Dreamina also demonstrate strong performance, with fewer than 1% of their videos missing any body parts.
+In contrast, MagicTime and LTX perform poorly, with 16.8% and 22.3% of their videos lacking body parts, respectively.
+These results indicate that these models struggle with maintaining consistency between text prompts and visual content.
+In terms of distortions, Sora, Wan, and Kling show excellent results, particularly Sora, which has only 0.75% of videos with noticeable facial distortions.
+Ying and Dreamina also exhibit good performance with low distortion rates.
+However, StableVideo, AnimateDiff, Show1, and Gen-2 show moderate results with some distortions.
+Open-Sora and MagicTime demonstrate significant distortion, with certain body parts exceeding 95%.
+This suggests these models struggle with maintaining consistency between text descriptions and the generated visuals.
+B.4. The Relationship between Semantic Distortion Identification and Visual Quality Scoring
+Figure 13. The correlation between the semantic distortion identification and subjective quality.
 We calculated the correlation between the three subjective quality scores and the distortion identification labels, as shown in Figure 13.
@@
 For example, in assessing the correlation between facial distortion and quality scores, we considered only videos where a face was visible.
-8.4.1 Body Parts Distortion
+B.4.1. Body Parts Distortion
 Overall, the quality of T2V models across the three evaluated dimensions demonstrates a certain degree of correlation with the distortions in the five body parts.
-Among these, facial distortion has the most significant impact on human appearance quality, followed by distortions in the torso and legs.
+Among these, facial distortion has the most significant impact on human appearance quality.
 This pattern aligns with the natural human perception process, which prioritizes facial recognition before focusing on the torso and other body parts.
 Of the three quality dimensions, HA Quality exhibits the strongest correlation with body part distortions, followed by AC Quality, with OV Quality showing the weakest correlation.
-This suggests that body part distortions most directly influence perceptions of human appearance quality, with a lesser effect on the continuity of human actions, and the least impact on overall video quality.
-Interestingly, for certain models, such as VideoCrafter2, quality scores show a negative correlation with distortions in the arms and feet.
-This implies that viewers’ attention is drawn more to prominent areas like the face and torso, making distortions in less visually critical body parts less influential.
-8.4.2 Body Parts Presence
+This suggests that body part distortions most directly influence perceptions of human appearance quality, with a lesser effect on the continuity of human actions and overall video quality.
+Interestingly, for certain models, such as VideoCrafter2, quality scores show a negative correlation with distortions in the hands and feet.
+This implies that viewers’ attention is drawn more to prominent areas like the face and other body parts, making distortions in less visually critical body parts less influential.
+B.4.2. Body Parts Presence
 We also calculated the correlation between the presence of five body parts and the subjective quality scores.
@@
 Interestingly, the presence of legs and feet often correlates negatively with all three quality dimensions, suggesting that their inclusion tends to degrade overall video quality.
-This underscores the current limitations of T2V models in accurately generating human legs and feet, as distortions in these areas significantly impact viewers’ subjective perception of video quality across all dimensions.
-9 Experiment Settings
-9.1 Evaluation Metrics
+This underscores the current limitations of T2V models in accurately generating human legs and feet, as distortions in these areas significantly impact viewers’ subjective perception of video quality across three dimensions.
+Appendix C Experiment Settings
+C.1. Evaluation Metrics
 Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC) are widely utilized metrics to quantify the relationship between predicted scores and ground truth scores.
@@
 It is mathematically defined as:
-$SRCC=1-\frac{6\sum_{n=1}^{N}(v_{n}-p_{n})^{2}}{N(N^{2}-1)},$ (12)
+(12) $SRCC=1-\frac{6\sum_{n=1}^{N}(v_{n}-p_{n})^{2}}{N(N^{2}-1)},$
 where $N$ denotes the number of data points, $v_{n}$ represents the rank of the $n$-th ground truth score, and $p_{n}$ represents the rank of the corresponding predicted score.
@@
 It is expressed as:
-$PLCC=\frac{\sum_{n=1}^{N}(y_{n}-\bar{y})(\hat{y}_{n}-\bar{\hat{y}})}{\sqrt{\sum_{n=1}^{N}(y_{n}-\bar{y})^{2}\sum_{n=1}^{N}(\hat{y}_{n}-\bar{\hat{y}})^{2}}},$ (13)
+(13) $PLCC=\frac{\sum_{n=1}^{N}(y_{n}-\bar{y})(\hat{y}_{n}-\bar{\hat{y}})}{\sqrt{\sum_{n=1}^{N}(y_{n}-\bar{y})^{2}\sum_{n=1}^{N}(\hat{y}_{n}-\bar{\hat{y}})^{2}}},$
 where $y_{n}$ and $\hat{y}_{n}$ denote the $n$-th ground truth and predicted scores, respectively, and $\bar{y}$ and $\bar{\hat{y}}$ are their corresponding means.
@@
 Unlike SRCC, which evaluates relative rankings, PLCC assesses the accuracy of predicted scores with respect to their numerical values, making it a critical metric for evaluating absolute predictive precision.
-9.2 Compared Quality Metrics
-9.2.1 Image Quality Assessment Methods
+C.2. Details of Evaluation for Distortion Identification
+In this study, we obtain outputs from Multi-Modal Large Language Models (MLLMs) through API calls or local deployment.
+We begin by extracting up to $32$ frames from each video, or all frames if fewer than $32$ are available.
+These frames, together with a prompt designed for the Distortion Identification task, are then fed into the MLLM.
+Finally, the raw outputs from the model undergo a post-processing procedure to produce the final results.
+The detailed prompts are listed below.
+For body part occurrences:
+# Suppose you are an expert in video content analysis.
+Please carefully examine and analyze these video frames, and determine whether a person’s [body part] appears in this video.
+Answer with ”Yes” or ”No”.
+#.
+For body part distortions:
+# Suppose you are an expert in video content analysis.
+Sequentially assemble the provided frames to create a complete video.
+Then meticulously examine and analyze the video, and determine whether there is any incompleteness, unrealistic appearance, or discontinuous movements of the person’s [body part] in this video.
+Answer with ”Yes” or ”No”.
+#.
+C.3. Compared Quality Metrics
+C.3.1. Image Quality Assessment Methods
 In all the IQA methods mentioned below, we uniformly extract 8 frames from each video in the Human-AGVQA dataset.
 These frames are then fed into the corresponding IQA methods to obtain 8 quality scores, and the average of these scores is taken as the method’s output.
-NIQE [58] is a no-reference image quality assessment (IQA) model that evaluates the quality of images based on natural scene statistics (NSS) without relying on human-rated training data.
+NIQE (Mittal et al., 2012b) is a no-reference image quality assessment (IQA) model that evaluates the quality of images based on natural scene statistics (NSS) without relying on human-rated training data.
 For the Human-AGVQA dataset, we applied NIQE to assess image quality using selected video frames.
-BRISQUE [57] leverages directional contrast features extracted from the Curvelet domain to assess the quality of contrast-distorted images.
+BRISQUE (Mittal et al., 2012a) leverages directional contrast features extracted from the Curvelet domain to assess the quality of contrast-distorted images.
 For the Human-AGVQA dataset, we utilized this approach to evaluate image quality based on its directional contrast features.
-HyperIQA [68] introduces a self-adaptive hyper network for blind image quality assessment (BIQA) in the wild, addressing diverse content and distortion types in real-world images.
+HyperIQA (Su et al., 2020) introduces a self-adaptive hyper network for blind image quality assessment (BIQA) in the wild, addressing diverse content and distortion types in real-world images.
 For the Human-AGVQA dataset, we applied HyperIQA to assess image quality by adaptively learning perception rules and predicting quality scores.
-UNIQUE [101] is a unified uncertainty-aware BIQA model designed to handle synthetic and realistic distortions.
+UNIQUE (Zhang et al., 2021) is a unified uncertainty-aware BIQA model designed to handle synthetic and realistic distortions.
 The model predicts both quality scores and associated uncertainties, enabling a probabilistic interpretation of image quality.
 For the Human-AGVQA dataset, we used UNIQUE to assess video frame quality based on its learning-to-rank framework.
-MUSIQ [34] introduces a multi-scale image quality Transformer designed to handle full-resolution images with varying sizes and aspect ratios.
+MUSIQ (Ke et al., 2021) introduces a multi-scale image quality Transformer designed to handle full-resolution images with varying sizes and aspect ratios.
 The model uses hash-based 2D spatial embedding and scale embedding to capture positional and scale information effectively.
 For the Human-AGVQA dataset, we applied MUSIQ to evaluate video frame quality by leveraging its multi-scale feature extraction capability.
-StairIQA [72] introduces a staircase network that integrates features from multiple convolutional layers of a CNN to capture better both low-level visual details and high-level semantic information for in-the-wild image quality assessment.
+StairIQA (Sun et al., 2023) introduces a staircase network that integrates features from multiple convolutional layers of a CNN to capture better both low-level visual details and high-level semantic information for in-the-wild image quality assessment.
 For the Human-AGVQA dataset, we used the official implementation and adopted this approach to evaluate image quality without modifications.
-CLIP-IQA [81] is a visual perception assessment model that utilizes the pre-trained CLIP framework to evaluate both the quality and abstract attributes of images.
+CLIP-IQA (Wang et al., 2023a) is a visual perception assessment model that utilizes the pre-trained CLIP framework to evaluate both the quality and abstract attributes of images.
 It introduces an antonym prompt pairing strategy (e.g., Good photo vs.
@@
 For the Human-AGVQA dataset, we used the official implementation, applying CLIP-IQA to assess video frame quality.
-LIQE [102] is a multitask learning framework for blind image quality assessment (BIQA) that leverages vision-language correspondence.
+LIQE (Zhang et al., 2023b) is a multitask learning framework for blind image quality assessment (BIQA) that leverages vision-language correspondence.
 LIQE uses CLIP for feature embeddings and optimizes tasks jointly by marginalizing probabilities over a textual template that describes scene, distortion, and quality.
 For the Human-AGVQA dataset, we used the official implementation, applying the quality assessment textual template to evaluate the video frame quality.
-MA-AGIQA [84] introduces a framework for AI-generated image quality assessment by combining large multi-modality models (LMMs) with traditional deep neural networks (DNNs).
+MA-AGIQA (Wang et al., 2024b) introduces a framework for AI-generated image quality assessment by combining large multi-modality models (LMMs) with traditional deep neural networks (DNNs).
 For the Human-AGVQA dataset, we used the official implementation, applying MA-AGIQA to evaluate the video frame quality without modifications.
-9.2.2 Video Quality Assessment Methods
-TLVQM [38] is a two-level video quality model that first computes low-complexity features across the entire video and then selects representative frames based on these features to extract high-complexity features.
-RAPIQUE [78] is a video quality assessment model that combines spatial-temporal scene statistics with high-level features extracted using deep CNNs.
+C.3.2. Video Quality Assessment Methods
+TLVQM (Korhonen, 2019) is a two-level video quality model that first computes low-complexity features across the entire video and then selects representative frames based on these features to extract high-complexity features.
+RAPIQUE (Tu et al., 2021b) is a video quality assessment model that combines spatial-temporal scene statistics with high-level features extracted using deep CNNs.
 For the Human-AGVQA dataset, we used the official implementation, applying sparse frame sampling and integrating spatial and temporal features to evaluate video quality.
-VIDEVAL [77] applies a feature selection strategy based on efficient blind VQA models.
+VIDEVAL (Tu et al., 2021a) applies a feature selection strategy based on efficient blind VQA models.
 Using the official open-source code, we converted Human-AGVQA videos from RGB to YUV420 format for feature extraction.
-Patch-VQ [96] is a video quality model that extracts spatial and temporal features from video frames and patches to analyze the relationship between local distortions and overall video quality.
+Patch-VQ (Ying et al., 2021) is a video quality model that extracts spatial and temporal features from video frames and patches to analyze the relationship between local distortions and overall video quality.
 For the Human-AGVQA dataset, we used Patch-VQ’s official implementation, dividing each video into space-time patches for feature extraction and quality prediction.
-SimpleVOA [71] uses an end-to-end network to extract spatial features directly from video frames and motion features to capture temporal distortions.
+SimpleVOA (Sun et al., 2022) uses an end-to-end network to extract spatial features directly from video frames and motion features to capture temporal distortions.
 A pre-trained SlowFast model processes 8 uniformly sampled frames rescaled to a height of 520.
 We used the official SimpleVOA model and finetuned it on Human-AGVQA.
-FAST-VQA [87] introduces a grid mini-patch sampling (GMS) strategy to balance local quality assessment with global context while reducing computational costs for high-resolution videos.
+FAST-VQA (Wu et al., 2022) introduces a grid mini-patch sampling (GMS) strategy to balance local quality assessment with global context while reducing computational costs for high-resolution videos.
 We used the official FAST-VQA-B model and finetuned it on Human-AGVQA.
-DOVER [88] is a video quality evaluator that assesses quality from technical and aesthetic perspectives. We used the official code without modifications.
-T2VQA [39] is a transformer-based model designed to evaluate text-to-video quality, focusing on text-video alignment and video fidelity.
+DOVER (Wu et al., 2023b) is a video quality evaluator that assesses quality from technical and aesthetic perspectives. We used the official code without modifications.
+T2VQA (Kou et al., 2024) is a transformer-based model designed to evaluate text-to-video quality, focusing on text-video alignment and video fidelity.
 It uses BLIP and Swin-T for feature extraction, followed by feature fusion through a cross-attention module and quality regression using a large language model.
 On the Human-AGVQA dataset, we used the official T2VQA model and finetuned it on Human-AGVQA.
-Table 7: Details of the model structure for the proposed model.
+Table 7. Details of the model structure for the proposed model.
 Training Hyper-Parameters Name/Value More Information
@@
 Text Feature Extraction PickScore backbone: CLIP-ViT-H/14
-UGVQ [103] is a unified framework for video quality assessment, focusing on spatial quality, temporal coherence, and text-to-video alignment.
+UGVQ (Zhang et al., 2024) is a unified framework for video quality assessment, focusing on spatial quality, temporal coherence, and text-to-video alignment.
 It utilizes features from CLIP for text and visual representations and SlowFast for motion representation.
 For the Human-AGVQA dataset, we employed UGVQ’s feature extraction and fusion modules to evaluate video quality comprehensively across these three dimensions.
-EvalCrafter [54] is an evaluation framework for text-to-video (T2V) generative models, focusing on video quality, text-video alignment, motion quality, and temporal consistency.
+EvalCrafter (Liu et al., 2024a) is an evaluation framework for text-to-video (T2V) generative models, focusing on video quality, text-video alignment, motion quality, and temporal consistency.
 It uses a benchmark of 700 prompts and 17 evaluation metrics to analyze T2V model performance.
 For overall video quality, we directly used implementation code in EvalCrafter on the Human-AGVQA dataset without specific changes.
-Q-Align [89] introduces a methodology for teaching large multi-modality models (LMMs) to assess video quality using discrete text-defined levels such as excellent, good, and poor.
+Q-Align (Wu et al., 2023c) introduces a methodology for teaching large multi-modality models (LMMs) to assess video quality using discrete text-defined levels such as excellent, good, and poor.
 It converts mean opinion scores (MOS) into these rating levels for training and infers scores by weighting the predicted probabilities of the levels.
 For the Human-AGVQA dataset, we used the official code for inference without modifications.
-9.2.3 Action Quality Assessment Methods
-USDL [74] is an uncertainty-aware score distribution learning approach for action quality assessment (AQA), treating actions as instances associated with score distributions.
+C.3.3. Action Quality Assessment Methods
+USDL (Tang et al., 2020) is an uncertainty-aware score distribution learning approach for action quality assessment (AQA), treating actions as instances associated with score distributions.
 For the Human-AGVQA dataset, since all videos contain more than 16 frames, each video is evenly divided into ten segments.
@@
 These normalized scores are then used to construct Gaussian distributions with the normalized value as the mean.
-ACTION-NET [99] is a hybrid attention network tailored for action quality assessment (AQA) in long videos.
+ACTION-NET (Zeng et al., 2020) is a hybrid attention network tailored for action quality assessment (AQA) in long videos.
 It combines dynamic video information with static postures of action subjects in selected frames.
@@
 For the static stream, the first, middle, and last frames are sampled, and the action subjects are cropped using the original detection algorithm.
-Table 8: Details of the hyper-parameters for the model training.
+Table 8. Details of the hyper-parameters for the model training.
 Training Hyper-Parameters Name/Value
@@
 optimizer Adam
-CoRe [97] evaluates action quality by regressing relative scores against a reference video with similar attributes, such as action category.
+CoRe (Yu et al., 2021) evaluates action quality by regressing relative scores against a reference video with similar attributes, such as action category.
 Due to differences in categorization strategies between the Human-AGVQA and AQA-7 datasets, we randomly selected an exemplar video generated by another T2V model within the same prompt category.
 Each video was divided into four equal segments, each containing four consecutive frames.
-TSA [92] is a temporal segmentation attention module designed for procedure-aware cross-attention learning after spatial-temporal feature extraction.
+TSA (Xu et al., 2022) is a temporal segmentation attention module designed for procedure-aware cross-attention learning after spatial-temporal feature extraction.
 Similar to CoRe, each video is evenly divided into segments of four consecutive frames and processed using the I3D model.
-For Motion Smoothness, Temporal Flickering, Action-Score, and Flow-Score metrics, we directly used their respective implementation code in VBench [33] and EvalCrafter [54] without specific changes.
-9.3 Training details
+For Motion Smoothness, Temporal Flickering, Action-Score, and Flow-Score metrics, we directly used their respective implementation code in VBench (Huang et al., 2024) and EvalCrafter (Liu et al., 2024a) without specific changes.
+C.4. Training details
 The detailed structure of the proposed GHVQ metric is provided in Table 7. The corresponding hyperparameters for model training are listed in Table 8.
-9.4 Ablation Study: The Effectiveness of Different Backbones
+C.5. Ablation Study: The Effectiveness of Different Backbones
 To further validate the effectiveness of the selected backbone networks, we conducted comparative experiments on four extraction modules: body-part segmentation network, holistic quality feature extraction network, text feature extraction network, and action quality analyzer.
-9.4.1 Body-part Segmentation Network
+C.5.1. Body-part Segmentation Network
 For the body-part segmentation network, performance differences between methods are more pronounced, as shown in Table 9.
@@
 Sapiens’ strong performance is also reflected in distortion identification.
-Table 9: Ablation Study of Body Segmentation Network.
+Table 9. Ablation Study of Body Segmentation Network.
 Methods HA Quality AC Quality OV Quality Distortion Identification
 SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
-None 0.659 0.671 0.692 0.704 0.745 0.752 0.581 0.524
-OpenPose [4] 0.757 0.764 0.740 0.748 0.753 0.751 0.745 0.667
-Deeplabv3+ [10] 0.712 0.725 0.724 0.736 0.758 0.767 0.649 0.546
-JPPNet [50] 0.685 0.692 0.726 0.732 0.773 0.782 0.582 0.475
-HRNet [69] 0.772 0.783 0.704 0.714 0.763 0.773 0.675 0.588
-Sapiens [36] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
-Table 10: Ablation Study of Text Feature Extraction Network.
+None 0.651 0.666 0.685 0.701 0.742 0.746 59.31 53.22
+OpenPose (Cao et al., 2017) 0.749 0.759 0.738 0.744 0.745 0.747 75.58 67.74
+Deeplabv3+ (Chen et al., 2018) 0.706 0.718 0.719 0.731 0.754 0.760 66.33 56.05
+JPPNet (Liang et al., 2018) 0.677 0.686 0.722 0.726 0.759 0.768 59.55 48.68
+HRNet (Sun et al., 2019) 0.767 0.775 0.699 0.706 0.757 0.764 68.91 59.94
+Sapiens (Khirodkar et al., 2024) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
+Table 10. Ablation Study of Text Feature Extraction Network.
 Methods HA Quality AC Quality OV Quality Distortion Identification
 SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
-None 0.781 0.789 0.754 0.765 0.773 0.776 0.802 0.736
-CLIPScore [27] 0.795 0.803 0.758 0.764 0.778 0.778 0.810 0.737
-BLIP [45] 0.806 0.812 0.772 0.782 0.768 0.773 0.810 0.744
-viCLIP [85] 0.803 0.807 0.769 0.773 0.767 0.774 0.813 0.739
-ImageReward [93] 0.801 0.805 0.775 0.778 0.772 0.779 0.811 0.735
-HPSv2 [91] 0.810 0.817 0.779 0.790 0.780 0.787 0.814 0.751
-PickScore [37] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
-9.4.2 Holistic Quality Feature Extraction Network
+None 0.774 0.779 0.737 0.742 0.759 0.760 80.53 74.16
+CLIPScore (Hessel et al., 2021) 0.791 0.795 0.741 0.748 0.763 0.765 81.28 74.11
+BLIP (Li et al., 2022a) 0.798 0.805 0.760 0.765 0.754 0.755 81.43 74.75
+viCLIP (Wang et al., 2022) 0.799 0.799 0.754 0.761 0.751 0.758 81.71 74.23
+ImageReward (Xu et al., 2023) 0.797 0.798 0.763 0.768 0.759 0.764 81.66 74.09
+HPSv2 (Wu et al., 2023a) 0.802 0.808 0.762 0.771 0.765 0.771 81.62 75.23
+PickScore (Kirstain et al., 2023) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
+C.5.2. Holistic Quality Feature Extraction Network
 For holistic quality features, PickScore achieves the highest scores across all quality metrics, as shown in Table 11.
@@
 It is worth noting that the performance of HPS v2 and CLIP-H/14 (LAION) is very close to that of Pickscore.
-Table 11: Ablation Study of Holistic Quality Feature.
+Table 11. Ablation Study of Holistic Quality Feature.
 Methods HA Quality AC Quality OV Quality Distortion Identification
 SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
-None 0.704 0.722 0.681 0.695 0.727 0.734 0.761 0.710
-CLIPScore [27] 0.728 0.739 0.712 0.719 0.743 0.748 0.772 0.717
-viCLIP [85] 0.723 0.731 0.723 0.733 0.732 0.744 0.789 0.731
-CLIP-B/16 (OpenAI) [64] 0.726 0.735 0.717 0.728 0.738 0.749 0.792 0.735
-CLIP-B/32 (OpenAI) [64] 0.739 0.745 0.731 0.742 0.746 0.752 0.787 0.729
-CLIP-L/14 (OpenAI) [64] 0.752 0.757 0.743 0.749 0.751 0.766 0.798 0.740
-CLIP-L/14 (LAION) [12] 0.776 0.785 0.768 0.777 0.760 0.768 0.802 0.743
-CLIP-H/14 (LAION) [12] 0.799 0.812 0.785 0.791 0.772 0.780 0.811 0.745
-CLIP-G/14 (LAION) [12] 0.795 0.805 0.782 0.786 0.768 0.778 0.805 0.737
-HPSv2 [91] 0.809 0.814 0.788 0.790 0.779 0.784 0.814 0.750
-PickScore [37] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
-Table 12: Ablation Study of Action Quality Analyzer.
+None 0.696 0.713 0.666 0.682 0.709 0.717 76.61 71.49
+CLIPScore (Hessel et al., 2021) 0.719 0.729 0.699 0.702 0.730 0.734 77.78 72.21
+viCLIP (Wang et al., 2022) 0.713 0.723 0.710 0.719 0.720 0.727 79.43 73.63
+CLIP-B/16 (OpenAI) (Radford et al., 2021b) 0.718 0.726 0.703 0.712 0.723 0.732 79.87 73.97
+CLIP-B/32 (OpenAI) (Radford et al., 2021b) 0.732 0.739 0.714 0.726 0.731 0.735 79.22 73.54
+CLIP-L/14 (OpenAI) (Radford et al., 2021b) 0.743 0.748 0.729 0.736 0.736 0.750 80.27 74.22
+CLIP-L/14 (LAION) (Cherti et al., 2023) 0.766 0.779 0.751 0.761 0.747 0.752 80.54 74.55
+CLIP-H/14 (LAION) (Cherti et al., 2023) 0.793 0.804 0.769 0.774 0.758 0.764 81.18 75.01
+CLIP-G/14 (LAION) (Cherti et al., 2023) 0.787 0.795 0.767 0.769 0.755 0.761 80.81 74.13
+HPSv2 (Wu et al., 2023a) 0.801 0.805 0.771 0.776 0.765 0.767 81.39 75.24
+PickScore (Kirstain et al., 2023) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
+Table 12. Ablation Study of Action Quality Analyzer.
 Methods HA Quality AC Quality OV Quality Distortion Identification
 SRCC PLCC SRCC PLCC SRCC PLCC Occ Dis
-None 0.767 0.770 0.747 0.748 0.728 0.740 0.753 0.704
-I3D [5] 0.781 0.789 0.752 0.762 0.709 0.713 0.811 0.739
-C3D [76] 0.797 0.798 0.760 0.773 0.709 0.713 0.811 0.739
-X3D [18] 0.808 0.812 0.784 0.790 0.764 0.769 0.813 0.747
-SlowFast [19] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
-9.4.3 Text Feature Extraction Network
+None 0.751 0.762 0.734 0.735 0.712 0.723 76.15 70.72
+I3D (Carreira and Zisserman, 2017b) 0.764 0.777 0.735 0.748 0.695 0.701 81.83 73.73
+C3D (Tran et al., 2015) 0.782 0.788 0.748 0.755 0.696 0.699 81.71 74.18
+X3D (Feichtenhofer, 2020) 0.793 0.801 0.767 0.774 0.750 0.757 81.96 74.84
+SlowFast (Feichtenhofer et al., 2019) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
+C.5.3. Text Feature Extraction Network
 For text feature extraction, PickScore achieves the highest scores across all quality metrics, as shown in Table 10.
@@
 In contrast, HPS v2 is trained on the more extensive HPD v2 dataset, characterized by its larger scale and greater diversity in both image content and preference distributions.
-9.4.4 Action Quality Analyzer
+C.5.4. Action Quality Analyzer
 For the action quality analyzer, the SlowFast method demonstrates superior performance across all quality metrics, particularly in HA Quality and AC Quality, as shown in Table 12.